### v1

# S1-Bench: A Simple Benchmark for Evaluating System 1 Thinking Capability of Large Reasoning Models

Abstract
We introduce S1-Bench, a novel benchmark designed to evaluate Large Reasoning Models’ (LRMs) performance on simple tasks that favor intuitive system 1 thinking rather than deliberative system 2 reasoning. While LRMs have achieved significant breakthroughs in complex reasoning tasks through explicit chains of thought, their reliance on deep analytical thinking may limit their system 1 thinking capabilities. Moreover, a lack of benchmark currently exists to evaluate LRMs’ performance in tasks that require such capabilities. To fill this gap, S1-Bench presents a set of simple, diverse, and naturally clear questions across multiple domains and languages, specifically designed to assess LRMs’ performance in such tasks. Our comprehensive evaluation of 22 LRMs reveals significant lower efficiency tendencies, with outputs averaging 15.5 times longer than those of traditional small LLMs. Additionally, LRMs often identify correct answers early but continue unnecessary deliberation, with some models even producing numerous errors. These findings highlight the rigid reasoning patterns of current LRMs and underscore the substantial development needed to achieve balanced dual-system thinking capabilities that can adapt appropriately to task complexity [^1] The code and benchmark can be found in https://github.com/WYRipple/S1_Bench..
[…]
tcb@breakable
S1-Bench: A Simple Benchmark for Evaluating System 1 Thinking Capability of Large Reasoning Models
Wenyuan Zhangthanks:  denotes equal contribution. denotes corresponding author., Shuaiyi Nie, Xinghua Zhang, Zefeng Zhang, Tingwen Liu Institute of Information Engineering, Chinese Academy of Sciences School of Cyber Security, University of Chinese Academy of Sciences {zhangwenyuan,nieshuaiyi,liutingwen}@iie.ac.cn
Figure 1: LRMs exhibit under-accuracy and overthinking on simple problems. Shapes represent organizations, colors represent base model families, with darker colors indicating larger models, and connecting lines represent the relationships between model families and training.
1 Introduction
“Simplicity is the ultimate sophistication.”
— Leonardo da Vinci
Recent advances in Large Reasoning Models (LRMs), notably OpenAI’s o1/o3 OpenAI (2024) and the DeepSeek-R1 Guo et al. (2025) series, have propelled the development of Large Language Models (LLMs) by explicitly generating external chain-of-thought (COT) Wei et al. (2022) before producing final answers, achieving remarkable performance on complex reasoning tasks. Unlike traditional LLMs that primarily rely on intuitive, heuristic system 1 thinking, LRMs enhance their performance by employing analytical, deliberative system 2 thinking Qu et al. (2025); Li et al. (2025b), which includes sophisticated strategies such as self-reflection, multi-path exploration Li et al. (2025a); Yeo et al. (2025). Despite the benefits of system 2 thinking, a rigid reliance on this cautious thinking mode can also introduce new limitations, such as performance degradation in specific contexts Feng et al. (2025); Zhao et al. (2025) or overthinking Chen et al. (2024); Kumar et al. (2025a).
However, a research gap exists in thoroughly investigating how LRMs’ over-reliance on system 2 thinking affects their performance when confronting extremely simple questions better suited for intuition-driven system 1 processing. Specifically, existing research primarily analyzes LRM behaviors in complex tasks that naturally align with deliberative system 2 thinking, such as MATH for math problems Luo et al. (2025); Yang et al. (2025); Chen et al. (2024) or IFEval for instruction following Zhao et al. (2025). Although some benchmarks involve relatively simple questions for human Miao et al. (2021); Yan et al. (2025); Chiang and Lee (2024), these tasks are often adversarial or non-trivial for LLMs, requiring moderate reasoning and can not be effectively solved through intuitive system 1 processing. Thus, a benchmark to assess the system 1 thinking capability of LRMs is still lacking, further hindering our understanding of LRMs’ cognitive flexibility between the two systems Ziabari et al. (2025); Qu et al. (2025).
To fill this research gap, we introduce the System 1 Thinking Capability Benchmark (S1-Bench), which measures the performance of LRMs across various simple tasks that commonly encountered in real-world applications. S1-Bench has the following three characteristics: (1) Simple. The questions are not hard for humans and can be easily answered by LLMs. LLMs with 7-9B parameters can robustly provide correct answers through direct responses when sampled across multiple temperatures. (2) Diverse. S1-Bench is not limited to simple reasoning problems; it encompasses four major categories and 28 subcategories in two languages (English and Chinese), including commonsense knowledge, loosely constrained instruction following, and straightforward analytical problems. (3) Natural. The questions are clear, without any misleading elements or ambiguities, ensuring they can be answered intuitively.
S1-Bench conducted extensive evaluations across 22 LRMs, yielding the following key findings: (1) Current LRMs exhibit inefficiency and lack system 1 thinking capabilities, with average output lengths 15.5 times longer than small LLMs. (2) LRMs often arrive at correct answers early in their thinking process but continue unnecessary reasoning with gradually increasing content similarity. (3) Despite employing deep reasoning, LRMs exhibit accuracy degradation on simple questions compared to traditional LLMs. (4) LRMs show the ability to prejudge the simplicity of questions, yet they still exhibit inefficiency. These findings emphasize the significant distance LRMs must traverse to become powerful dual-system compatible models.
Our contributions can be summarized as follows:
  • To the best of our knowledge, S1-Bench is the first benchmark to evaluate the system 1 thinking capabilities of LRMs.
  • We introduce a workflow for constructing simple dataset suitable for system 1 evaluation.
  • Extensive experiments reveal the inefficiency and under-accuracy of LRMs on simple questions that are suited for system 1 thinking.
2 Related work
Figure 2: Construction workflow for S1-Bench and an illustrative example from each major category.
2.1 Large Reasoning Models
Large Reasoning Models (LRMs), characterized by explicitly generating external thinking processes before final answers Kumar et al. (2025b); Chen et al. (2025), achieve a paradigm shift from intuitive system 1 thinking to deliberative system 2 reasoning compared to traditional LLMs Li et al. (2025b); Qu et al. (2025), thus achieving superior performance on complex tasks. The development of recent LRMs has largely followed two main approaches: large-scale reinforcement learning (RL) and model distillation. Models trained via large-scale RL Guo et al. (2025); Team (2025b); Team et al. (2025) leverage reward-based optimization to gradually incentivize deliberative reasoning. In contrast, distillation-based LRMs OpenAI (2024); Min et al. (2024); Team (2025a); Ye et al. (2025); Muennighoff et al. (2025) acquire such abilities by transferring structured reasoning patterns from advanced teacher models.
2.2 Limitations of LRMs
While LRMs have shown significant performance gains through deliberate reasoning, rigid adherence to this overly cautious thinking can introduce new limitations. On one hand, intermediate reasoning steps can cause excessive token generation and unnecessary solving attempts Chen et al. (2024); Hashemi et al. (2025); Kumar et al. (2025a). On the other hand, LRMs’ performance can drop in specific contexts like safety scenarios Jiang et al. (2025); Zhao et al. (2025) and role-playing Feng et al. (2025). However, prior studies mainly evaluated LRMs on complex tasks that more suited for deliberative system 2 thinking. Our work examines how deliberative reasoning impacts extremely simple problems better matched to intuition-driven system 1 processing.
3 S1-Bench
We introduce S1-Bench, a bilingual, multi-domain benchmark designed to evaluate system 1 thinking capability of LRM on extremely simple questions. These questions are easily solvable by traditional LLMs and not hard for humans. S1-Bench, which covers both English and Chinese, is organized into four major categories: reasoning (Rsn), knowledge (Kno), instruction following (If) and analysis (Ana), representing major dimensions commonly employed in LLM capability evaluation Zheng et al. (2023); Chang et al. (2024).
This section begins with how simplicity is ensured, then the detailed construction workflow for S1-Bench, and concludes with an overview of the dataset statistics. Figure 2 shows the construction workflow and an illustrative example per category.
3.1 How to Ensure Simplicity?
We ensure questions are simple and suitable for system 1 thinking through two aspects.
3.1.1 A Priori Simplicity Constraints
We require all questions and answers to adhere to the following guidelines:
(1) Questions must be naturally, clearly expressed, and unambiguous, free of intentional traps.
(2) Answers must be unique or easily falsifiable (such as providing a three-letter English word).
We’ve further implemented category-specific constraints to ensure simplicity: Rsn: Limited to problems solvable with minimal reasoning or intuition. Kno: Restricted to common knowledge with unique, verifiable answers from sources like Wikipedia. If: Involve straightforward instructions without strict formatting requirements. Ana: Limited to questions whose answers can be directly inferred from the prompt, such as binary classification. These constraints ensure all questions remain straightforward for human respondents.
3.1.2 A Posteriori Simplicity Verification
Due to the biases existing between language models and humans Gallegos et al. (2024), questions that are simple for humans may be difficult for LLMs. Therefore, we introduce additional posteriori verification to ensure that questions are simple enough to be correctly and robustly answered by smaller LLMs from different families. This alleviates cognitive biases in traditional LLMs for system 1 thinking problems Li et al. (2025b).
3.2 Construction Workflow
To implement a priori constraints, we introduce data generators [^2] We select Claude-3.7-Sonnet and Qwen2.5-72B-Instruct. and quality discriminators [^3] We select GPT-4o and DeepSeek-V3-241226. that adhere to the a priori settings. To ensure posterior verification, we introduce simplicity validators [^4] We select four small LLMs: Qwen2.5-7B, Llama3.1-8B, Mistral-8B, and Gemma2-9B. The full model IDs are detailed in Table 4. and correctness evaluator [^5] We select GPT-4o.. The generators, discriminators, and evaluator are uniformly configured with temperature 0.7 and top-p=0.9, while validators are set with temperatures of 0.0, 0.2, and 0.4, each sampling 10 times. Additionally, we recruit three experienced graduate students as annotators, familiar with LLMs and fully comprehending the construction objectives of S1-Bench. For the prompts used in the workflow, refer to Table 11.
Subcategory Preparation. Initially, to ensure diversity, we draw on existing benchmark classification methods (e.g., MMLU, IFEval and GSM8K) and evaluation surveys Chang et al. (2024), selecting or merging existing subcategories, and designing new ones. We ensure that all subcategories meet the requirements for simplicity and provide an example question for each.
Generation and Modification. We use two data generators to create initial questions and answers for each candidate subcategory. The generation process adheres to prompt requirements (refer to Table 11), ensuring bilingual content that aligns with subcategory definitions while maintaining diverse perspectives. Each question is accompanied by auxiliary answers, with 50 initial bilingual question-answer pairs generated per subcategory. These pairs are independently assessed by three annotators and two quality discriminators , who ensure that the questions are unambiguous, have clear boundaries, and admit only one correct answer. For each question-answer pair, five independent assessments are conducted. Based on these evaluations, annotators discuss and collectively decide whether to retain, modify, or discard each question, ensuring alignment with the requirements outlined in Section 3.1.1. Retained questions are submitted to validators for testing, and an evaluator assesses the accuracy of responses. Questions with correct answers are included in S1-Bench, while others undergo reprocessing.
Questions with errors in any of the 30 sampling iterations undergo an iterative difficulty reduction process. These questions return to generators with specific prompts to decrease complexity, followed by the same discrimination, discussion, and evaluation procedures. Questions failing to achieve error-free performance after three difficulty reduction iterations are excluded from the workflow. The final S1-Bench comprises questions that satisfy both human-established a priori simplicity constraints and a posteriori simplicity verification.
Figure 3: Statistical distribution of token counts for S1-Bench questions.
Model t=0.0 t=0.2 t=0.4 Tokens
Gemma2-9B 100.00 100.00 100.00 38.77
Llama3.1-8B 100.00 100.00 100.00 42.00
Mistral-8B 100.00 100.00 100.00 44.38
Qwen2.5-7B 100.00 100.00 100.00 42.81
Qwen2.5-14B 99.74 99.76 99.76 40.00
Qwen2.5-32B 99.98 99.98 99.98 43.17
Qwen2.5-72B 100.00 100.00 100.00 44.61
Llama3.3-70B 100.00 99.76 99.76 53.71
DeepSeek-v3 100.00 100.00 100.00 79.53
Table 1: Average accuracy and response token count of different LLMs, each sampled 10 times at three temperature settings.
3.3 Benchmark Statistics
S1-Bench comprises 422 question-answer pairs across four major categories and 28 subcategories, balanced with 220 English and 202 Chinese questions. We present the category distribution and descriptions with examples for all subcategories, which can be found in Figure 8 and Table 15,16 in the appendix. Figure 3 shows the token length distribution, with questions averaging 14.46 tokens.
To validate the absence of response bias from validators, we tested S1-Bench on five LLMs that serve as base models for commonly used LRMs. As shown in Table 1, these models achieved an accuracy exceeding 99% under three temperature settings, confirming the simplicity and robustness of S1-Bench.
Model ID Size Loose Format Strict Format L-Corr $\uparrow$ S-Corr $\uparrow$ $\mathtt{ART}\downarrow$
pass@1$\uparrow$ acc@$\uparrow$ pass@1$\uparrow$ acc@$\uparrow$
Validator LLMs 7-9B 100.00 100.00 100.00 100.00 42.00
QwQ-32B 32B 100.00 100.00 100.00 100.00 100.00 100.00 720.10
Hunyuan-T1 99.91 99.53 99.91 99.53 100.00 100.00 542.31
Sky-T1-32B 32B 98.82 94.79 94.88 79.62 99.48 95.26 163.00
DS-R1 671B 100.00 100.00 100.00 100.00 100.00 100.00 646.40
DS-R1-70B 70B 99.48 97.39 99.38 96.92 100.00 99.91 453.81
DS-R1-32B 32B 99.72 98.82 99.72 98.82 100.00 100.00 429.91
DS-R1-14B 14B 99.57 97.87 99.57 97.87 100.00 100.00 475.46
DS-R1-8B 8B 97.44 97.16 97.39 97.16 99.76 99.53 452.11
DS-R1-7B 7B 95.21 85.78 95.21 85.78 99.24 99.24 454.55
DS-R1-1.5B 1.5B 81.47 54.50 81.47 54.50 97.58 97.58 489.54
Nemotron-49B 49B 99.15 97.39 99.15 97.39 100.00 100.00 362.54
Nemotron-8B 8B 86.16 69.91 79.81 59.00 99.43 84.31 372.57
L-R1-32B 32B 97.87 91.00 94.74 79.62 98.91 95.07 1095.36
L-R1-32B-DS 32B 99.57 98.10 99.57 98.10 99.81 99.81 524.12
L-R1-14B-DS 14B 99.05 95.97 99.05 95.97 99.19 99.19 693.19
L-R1-7B-DS 7B 94.64 83.65 94.64 83.65 99.76 99.67 496.47
s1.1-32B 32B 99.53 98.34 99.48 98.10 99.57 99.53 998.00
s1.1-14B 14B 97.63 93.60 97.25 91.94 97.77 97.39 839.86
s1.1-7B 7B 96.68 88.39 88.58 63.98 97.11 88.96 711.49
EXAONE-32B 32B 97.06 94.08 97.06 94.08 99.81 99.81 800.56
EXAONE-7.8B 7.8B 88.15 75.12 87.82 74.41 98.72 98.06 1046.87
EXAONE-2.4B 2.4B 72.42 56.16 72.32 56.16 97.44 97.25 1593.96
Table 2: Main results in the top-p sampling setting on the S1-Bench, sorted by model family. Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst.
4 Main Experiment
4.1 Baseline Models and Configurations
We evaluated 22 different LRMs, which were explicitly trained to first respond with a thinking process, and then generate a final answer. These LRMs include open-source families, such as DeepSeek Guo et al. (2025), Qwen Team (2025b), Nemotron Chris Alexiuk and Patel (2025), Light-R1 Wen et al. (2025), s1.1 Muennighoff et al. (2025), EXAONE Research et al. (2025), and Sky-T1 Griggs et al. (2025), as well as closed-source Hunyuan-T1 Tencent (2025), spanning from tiny (1.5B) to large (671B) parameter sizes [^6] More details in Appendix Table 5 (includes training data scale and algorithms).. Notably, OpenAI’s o-series models are not included as they do not disclose thinking processes to users. For each model, we consider two sets of generation configurations: Greedy sampling with temperature t = 0; Top-p sampling with temperature t = 0.6, top-p = 0.95 and sampling size k=5.
4.2 Evaluation Metrics
Format Metrics.
Generally, the output of LRMs should be separated by an “end thinking marker” (e.g., </think>) that distinguishes the thinking process from the final answer. However, LRMs do not always generate responses in correct format. To assess this, we calculate the percentage of responses that meet the required formatting criteria, categorized as follows (for top-p sampling setting, each metric is averaged across 5 generations):
S-Corr (Strict Format Correctness Rate): The percentage of responses that strictly adhere to the format, meaning they only contain one "end thinking marker" followed by a non-empty final answer.
L-Corr (Loose Format Correctness Rate): The percentage of responses that loosely adhere to the format, meaning they meet the format requirements except for responses with endless repetitions [^7] Detailed format analyses are in the Appendix C..
Model ID Size S1-Bench-$\mathtt{EN}$ S1-Bench-$\mathtt{ZH}$ Avg
Rsn Kno If Ana Avg Rsn Kno If Ana Avg
Gemma2-9B 9B 74.8 29.4 5.3 52.4 45.9 51.6 19.8 7.5 35.1 31.0 38.8
Llama3.1-8B 8B 91.0 35.4 12.4 61.9 56.0 44.0 28.3 15.2 18.7 26.7 42.0
Qwen2.5-7B 7B 65.5 46.3 6.4 49.6 46.5 50.5 46.6 9.8 36.9 38.8 42.8
Mistral-8B 8B 67.2 55.5 8.6 50.1 49.6 47.3 56.1 14.8 29.7 38.7 44.4
Column Avg - 74.6 41.6 8.2 53.5 49.5 48.3 37.7 11.8 30.1 33.8 42.0
Sky-T1-32B 32B 215.8 174.1 98.5 233.3 194.3 125.5 125.3 99.4 145.5 128.9 163.0
Nemotron-49B 49B 599.7 587.6 396.5 526.1 540.4 232.9 157.3 235.5 107.8 168.8 362.5
Nemotron-8B 8B 561.0 585.1 458.0 303.1 462.6 369.5 326.0 288.1 166.7 273.5 372.6
DS-R1-32B 32B 421.8 504.4 414.7 521.1 473.7 362.2 385.6 343.1 408.8 382.2 429.9
DS-R1-8B 8B 472.2 528.9 530.7 462.7 491.2 521.9 404.4 266.2 395.5 409.4 452.1
DS-R1-70B 70B 464.1 501.3 378.5 536.1 484.0 450.8 450.2 328.4 416.7 420.9 453.8
DS-R1-7B 7B 447.5 623.9 353.8 510.0 495.5 446.5 463.2 339.5 373.0 409.4 454.5
DS-R1-14B 14B 503.7 674.7 367.3 494.2 519.0 452.0 465.4 375.3 405.8 428.0 475.5
DS-R1-1.5B 1.5B 480.8 584.7 417.4 577.2 529.1 493.0 497.4 329.8 423.1 446.0 489.5
L-R1-7B-DS 7B 568.1 667.1 501.7 566.3 580.3 444.8 454.6 344.1 366.4 405.0 496.5
L-R1-32B-DS 32B 574.5 706.6 647.6 632.8 636.3 431.2 367.0 377.1 418.7 402.2 524.1
Hunyuan-T1 561.6 693.8 380.9 435.0 521.2 676.8 553.8 505.1 523.8 565.3 542.3
DS-R1 671B 786.1 723.8 711.4 529.2 672.5 727.3 638.5 607.9 533.9 617.9 646.4
L-R1-14B-DS 14B 951.0 1026.0 829.8 653.5 848.2 594.7 610.1 442.2 451.7 525.7 693.2
s1.1-7B 7B 1039.5 840.8 1923.2 529.4 929.9 489.6 351.3 1034.3 332.4 475.6 711.5
QwQ-32B 32B 873.3 808.1 520.8 634.7 722.4 866.9 707.3 613.3 667.7 717.6 720.1
EXAONE-32B 32B 1323.7 1057.6 1537.0 711.6 1086.4 703.2 348.6 1302.9 125.5 490.3 800.6
s1.1-14B 14B 871.8 746.2 2233.1 708.1 960.2 654.6 546.0 1512.6 579.7 710.7 839.9
s1.1-32B 32B 1077.9 889.7 2055.4 781.7 1081.7 995.6 765.2 1634.6 666.5 906.5 998.0
EXAONE-7.8B 7.8B 1498.3 1398.9 1775.7 882.4 1303.8 1410.3 497.8 1633.1 205.0 767.0 1046.9
L-R1-32B 32B 1614.0 1217.3 1996.9 930.1 1338.3 1035.6 737.7 1240.7 610.2 835.3 1095.4
EXAONE-2.4B 2.4B 1927.3 1426.2 1200.1 825.7 1320.7 2469.7 1622.6 2471.6 1511.2 1898.7 1594.0
Column Avg - 810.6 771.2 896.8 590.2 736.0 679.8 521.6 742.0 447.1 563.0 652.8
Improvement - $\times$10.9 $\times$18.5 $\times$109.7 $\times$11.0 $\times$14.9 $\times$14.1 $\times$13.8 $\times$62.7 $\times$14.8 $\times$16.7 $\times$15.5
Table 3: $\mathtt{ART}$ on the S1-bench across two languages and four main categories. Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst. Bold represents the maximum Improvement value for each language.
Efficiency Metrics.
We calculate the Average Response tokens (ART), which represent the average number of tokens in responses that meet the loose formatting requirements. Token counts are obtained using the Qwen2.5 tokenizer.
Accuracy Metrics.
We calculate accuracy metrics for both strict and loose formatting requirements, respectively. We use GPT-4o as the evaluator to assess the correctness of the responses [^8] If a final answer can be isolated, only the final answer is evaluated; otherwise, the entire response is assessed., with the evaluation prompt in Table 11. For greedy sampling, we directly calculate the accuracy rate, with complete results in Appendix D. For top-p sampling, we utilize two metrics:
Pass@1: Followed DeepSeek-R1 Guo et al. (2025), we calculate pass@1 to assess the percentage of correct responses among the k=5 generations. Specifically, it is defined as:
[…]
where $p_{i}$ is the correctness of the i-th generation.
Acc@k: Since S1-Bench is composed of extremely simple questions, we calculate acc@k. Specifically, acc@k=1 if all k responses are correct and acc@k = 0 otherwise. It is defined as:
[…]
Notably, S-Corr represents the upper bound for pass@1 and acc@5 under strict formatting requirements, with L-Corr following the same principle.
4.3 Main Results
Table 2 and Figure 1 presents the main results of LRMs on S1-Bench, revealing two key phenomena.
LRMs exhibit significantly lower efficiency than LLMs on S1-Bench, and no clear correlation is observed between ART and model size.
Notably, state-of-the-art LRMs, such as DeepSeek-R1 and QwQ-32B, do not demonstrate a distinct advantage in efficiency. In contrast, Sky-T1-32B, which undergoes specific optimizations to mitigate overthinking using SimPO, achieves the highest efficiency. The L-R1-DS 7B/14B/32B models generate longer responses than the DS-R1-7B/14B/32B models, and the former are further post-trained on the latter. This suggests that further post-training may improve reasoning ability for complex tasks, but at the cost of response efficiency for simple tasks. Finally, the s1.1 models generate considerably longer responses than the DeepSeek-R1-Distilled models. Despite both models being trained solely with SFT to acquire long-COT reasoning ability, the DeepSeek-R1-Distilled models use 800K training samples, while the s1.1 models are trained on only 1K. This discrepancy suggests that the smaller training set may lead to superficial imitation of long reasoning patterns, resulting in verbose thinking on simple questions.
Several LRMs exhibit under-accuracy on simple questions.
Our observations reveal that, despite employing deep reasoning, most LRMs tend to exhibit lower accuracy on simple questions compared to traditional LLMs. For example, DS-R1-1.5B and EXAONE-2.4B achieve just above 50% acc@k. Moreover, accuracy decreases with smaller model size. Finally, many LRMs struggle with robust correctness in top-p sampling, where acc@k is significantly lower than pass@1. This issue is particularly pronounced in smaller LRMs. For instance, DS-R1-1.5B achieved 81.47% pass@1 but only 54.50% acc@k.
5 Efficiency Analysis
Figure 4: (a) Comparison of Initial and Additional Thinking Costs for Each LRM. (b) Distribution of Solution Rounds for Each LRM.
5.1 ART Analysis of Question Types
To better understand the efficiency differences of LRMs across question types, we analyze $\mathtt{ART}$ across 4 main categories, 28 subcategories, and two languages, yielding the following insights.
LRMs exhibit the most significant increase in ART for instruction-following questions and tend to over-explore when the solution space is vast.
Although small LLMs generate the most concise responses for instruction-following questions across all main categories in S1-Bench, LRMs produce responses that are 109.7× longer in Chinese and 62.7× longer in English, making them the longest response type across all categories. We further analyze the subcategories of instruction-following questions. As shown in Figure 9, $\mathtt{ART}$ is notably longer in the subcategories of length constraints, character constraints, and sentence constraints. These three question types share a similar characteristic: their correctness is verifiable, but the solution space is vast. We find that, although the model quickly identifies a correct answer, it becomes trapped in the search space, continually exploring alternatives and failing to stop in time. This phenomenon is more pronounced in families with higher $\mathtt{ART}$, such as s1.1 and the EXAONE.
LRMs also demonstrate a significant increase in ART for reasoning tasks.
In Table 3, $\mathtt{ART}$ for reasoning questions also increases significantly, and Figure 9 further corroborates this phenomenon in $\mathtt{ART}$ in various reasoning subcategories. One possible explanation for this is that reasoning questions align closely with the post-training data distribution of LRMs, thereby further stimulating the long-chain reasoning behavior of LRMs.
LRMs consistently exhibit inefficiency in both Chinese and English.
In table 3, most LRMs show higher $\mathtt{ART}$ on English questions, while QwQ-32B and Hunyuan-T1 exhibit comparable $\mathtt{ART}$ across Chinese and English.
Figure 5: Maximum similarity between each segment and all preceding segments for LRMs across four categories.
5.2 Solution Analysis in Thinking Processes
To better understand the causes of inefficiency in LRMs on the S1-Bench, we analyze the thinking process where the final answer is correct and the format is strictly correct and non-empty [^9] We selected this subset because incorrect answers make it unclear whether LRMs are over-reasoning or under-reasoning, and malformed or empty thinking processes cannot be precisely extracted.. We begin by segmenting each thinking process into several solutions, each defined as a point at which LRMs explicitly arrive at a conclusion that directly aligns with the correct answer. The segmentation process is performed by DeepSeek-v3, with prompts detailed in Table 12. We then compute the average initial thinking cost for LRMs. For each sample, if the thinking process contains at least one solution, the cost is defined as the tokens counts in the first solution. If no clear and correct solution is provided, the cost is the total token counts in the thinking process. The reported value is the average across all selected samples. We find that:
The step-by-step reasoning in the initial thinking process is not the primary cause of the inefficiency in LRMs.
As shown in Figure 4 (a), although the total counts of thinking tokens varies considerably across LRMs, their initial thinking costs are similar and account for only a small fraction of the total.
Generating unnecessary solution rounds after reaching the correct answer is one of the reasons for the inefficiency of LRMs.
We further examine the distribution of solution rounds among various LRMs on the S1-Bench (Figure 4 (b)) and find that models with longer thinking processes tend to produce excessive solution rounds, repeatedly re-verifying simple problems that have already been solved. This redundant verification significantly contributes to computational inefficiency.
Figure 6: Distribution of the thinking process across four categories. FA and TP refer to Final Answer and Thinking Process, respectively. Green bars indicate cases where the final answer is correct, while red bars indicate cases where it is incorrect.
5.3 Redundancy in Thinking Processes
Information redundancy increases as reasoning sequences increase. We conduct a similarity analysis to analyze how information redundancy in the thinking processes of LRMs changes as reasoning sequences increase. Specifically, we first divide the complete thinking process into k equal-length segments [^10] We set k=15, changing its value does not affect the conclusions.. Then, we encode each segment using the all-MiniLM-L6-v2 model [^11] https://huggingface.co/sentence-transformers/allMiniLM-L6-v2. For each segment, we calculate the cosine similarity with all its preceding segments and use the maximum similarity as a measure of its information redundancy. As shown in Figure 5, information redundancy increases across all four main categories as reasoning sequences increase. Sky-T1-32B shows overall lower similarity, which stems from its shorter thinking process, but still demonstrates an upward trend.
6 Analysis of Errors
Figure 7 explores the relationship between thinking process quality and final answer accuracy. We focus on samples whose format is strictly correct and thinking process is non-empty. For samples with correct final answers, we categorize them based on whether the thinking process contains explicit incorrect conclusions in intermediate steps. For samples with incorrect final answers, we categorize them based on whether the correct answer is mentioned at least once during reasoning. We use DeepSeek-v3 for categorization, with the prompt provided in Table 13. Figure 6 presents the distribution of these four categories in different RLMs. Our analysis yields the following observations: (1) LRMs with lower accuracy often include incorrect intermediate conclusions in their reasoning, even when they ultimately reach correct final answers (light green). (2) Although LRMs sometimes reach the correct answer during reasoning, they may deviate and ultimately produce incorrect conclusions (light red).
7 Simplicity Prejudgement
Figure 7: Counting instances of prejudgment generated by LRMs for simple questions and $\mathtt{ART}$.
Finally, we discover an intriguing phenomenon: LRMs can prejudge certain simple questions. Specifically, we extract the first 50 tokens from each thinking process and use DeepSeek-v3 to identify sentences showing judgment, further extracting judgment words [^12] An example of prejudgment: “This question seems simple”. For prompt references, see Table 14.. Figure 7 displays the frequency of words judging questions as simple, with details provided in the Table 7. Our findings reveal:
LRMs possess the ability to prejudge question simplicity, especially in Chinese.
All LRMs exhibit prejudgment phenomena within their thinking processes, demonstrating an ability to directly assess question difficulty. Additionally, 14 LRMs exhibit significantly stronger prejudgment tendencies in Chinese, with this phenomenon concentrating primarily within the L-R1, DS-R1, QwQ, and Hunyuan model families.
Even with prejudgment, the thinking length of LRMs does not shorten.
As shown in Figure 7, the average $\mathtt{ART}$ of the thinking process exhibiting prejudgment does not decrease. We consider further exploration of this phenomenon as an important direction for future work.
These results suggest LRMs possess an inherent understanding of question difficulty. This opens a novel pathway toward dual-system compatibility for LRMs, which we identify as future work.
8 Conclusion
This paper presents S1-Bench, designed to evaluate LRMs’ performance on simple questions suited for system 1 thinking. Experiments reveal significant overthinking and performance degradation by LRMs on S1-Bench, highlighting the gap between current LRMs and dual-system compatibility.
Appendix A Subcategories in S1-Bench
Figure 8: S1-Bench Category Display. The inner circle represents four major categories, and the outer circle includes 28 subcategories.
Figure 8 shows the pie chart distribution of 28 subcategories in S1-Bench. For more details on the subcategories, please refer to Table 15,16.
Appendix B Baseline Model Details
Model Model ID URL
Qwen2.5-7B Qwen2.5-7B-Instruct https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
Llama3.1-8B Llama-3.1-8B-Instruct https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
Mistral-8B Ministral-8B-Instruct-2410 https://huggingface.co/mistralai/Ministral-8B-Instruct-2410
Gemma2-9B gemma-2-9b-it https://huggingface.co/google/gemma-2-9b-it
Qwen2.5-14B Qwen2.5-14B-Instruct https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
Qwen2.5-32B Qwen2.5-32B-Instruct https://huggingface.co/Qwen/Qwen2.5-32B-Instruct
Qwen2.5-72B Qwen2.5-72B-Instruct https://huggingface.co/Qwen/Qwen2.5-72B-Instruct
Llama3.3-70B Llama-3.3-70B-Instruct https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct
DeepSeek-v3 DeepSeek-V3-0324 https://huggingface.co/deepseek-ai/DeepSeek-V3-0324
Table 4: Mapping of LLM abbreviations and IDs used in this paper, with their open-source URLs.
Model ID Abbreviation Base Model Alg. Size
DeepSeek
DeepSeek-R1-Distill-Qwen-1.5B DS-R1-1.5B Qwen2.5-Math-1.5B SFT 800K
DeepSeek-R1-Distill-Qwen-7B DS-R1-7B Qwen2.5-Math-7B SFT 800K
DeepSeek-R1-Distill-Llama-8B DS-R1-8B Llama-3.1-8B SFT 800K
DeepSeek-R1-Distill-Qwen-14B DS-R1-14B Qwen2.5-14B SFT 800K
DeepSeek-R1-Distill-Qwen-32B DS-R1-32B Qwen2.5-32B SFT 800K
DeepSeek-R1-Distill-Llama-70B DS-R1-70B Llama-3.3-70B-Instruct SFT 800K
DeepSeek-R1 DS-R1 DeepSeek-V3-0324 SFT&RL 800K&–
Qwen
QwQ-32B QwQ-32B Qwen2.5-32B
qihoo360
Light-R1-7B-DS L-R1-7B-DS DeepSeek-R1-Distill-Qwen-7B SFT 3K
Light-R1-14B-DS L-R1-14B-DS DeepSeek-R1-Distill-Qwen-14B SFT&RL 3K&–
Light-R1-32B-DS L-R1-32B-DS DeepSeek-R1-Distill-Qwen-32B SFT 3K
Light-R1-32B L-R1-32B Qwen2.5-32B-Instruct SFT&DPO 73K&–
simplescaling
s1.1-7B s1.1-7B Qwen2.5-7B-Instruct SFT 1K
s1.1-14B s1.1-14B Qwen2.5-14B-Instruct SFT 1K
s1.1-32B s1.1-32B Qwen2.5-32B-Instruct SFT 1K
LG AI Research
EXAONE-Deep-2.4B EXAONE-2.4B EXAONE-3.5-2.4B-Instruct SFT&DPO&RL 1.6M&20K&10K
EXAONE-Deep-7.8B EXAONE-7.8B EXAONE-3.5-7.8B-Instruct SFT&DPO&RL 1.6M&20K&10K
EXAONE-Deep-32B EXAONE-32B EXAONE-3.5-32B-Instruct SFT&DPO&RL 1.6M&20K&10K
NVIDIA
Llama-3.1-Nemotron-Nano-8B-v1 Nemotron-8B Llama-3.1-8B-Instruct SFT&RL –&–
Llama-3.3-Nemotron-Super-49B-v1 Nemotron-49B Llama-3.3-70B-Instruct SFT&RL –&–
NovaSky
Sky-T1-32B-Flash Sky-T1-32B Qwen2.5-32B-Instruct SFT&SimPO 17K&10K
Table 5: The open-source LRMs details evaluated for S1-Bench.
Table 4 presents the abbreviations, IDs, and URLs of LLMs used in this paper. Table 5 displays the abbreviations, IDs, URLs, organizations, algorithms, and training data volumes of open-source LRMs evaluated in this study.
Appendix C Format Details
Format ID marker marker marker thinking final
(standard) (special) (number) process answer
Strict 100 $\surd$ 1 $\surd$ $\surd$
101 $\surd$ 1 $\times$ $\surd$
Loose (w/o Strict) 200 $\surd$ 1 $\surd$ $\times$
201 $\surd$ >1 $\surd$ $\surd$
202 $\surd$ >1 $\surd$ $\times$
203 $\surd$ >1 $\times$ $\surd$
204 $\times$ $\surd$ $\geq$1 $\surd$ $\surd$
205 $\times$ $\surd$ $\geq$1 $\surd$ $\times$
206 $\times$ $\surd$ $\geq$1 $\times$ $\surd$
207 $\times$ $\times$ 0 $\surd$
Error 300 $\surd$ $\surd$ $\geq$1 $\times$ $\times$
301 $\times$ $\times$ 0 $\infty$
Table 6: Format types.
Words Count
{CJK}UTF8gkai简单 466
straightforward 231
simple 118
{CJK}UTF8gkai不难 72
{CJK}UTF8gkai基础的 44
basic 24
{CJK}UTF8gkai直接 15
{CJK}UTF8gkai常见 10
{CJK}UTF8gkai基础 10
{CJK}UTF8gkai不复杂 7
simplest 7
pretty straightforward 5
easy 5
{CJK}UTF8gkai基本的 5
{CJK}UTF8gkai基本 3
Table 7: Top 15 phrases with prejudgment phenomenon, with counts summed across 22 LRMs.
Model Loose Format Error Format
Strict Format
100 101 200 201 202 203 204 205 206 207 300 301
QwQ-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
hunyuan-t1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Sky-T1-32B 95.26 0.00 0.62 0.09 0.19 0.00 0.28 0.00 0.00 3.03 0.00 0.52
DS-R1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-70B 99.91 0.00 0.09 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-14B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-8B 99.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24 0.00 0.24
DS-R1-7B 99.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.76
DS-R1-1.5B 97.58 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.42
Nemotron-49B 66.07 33.93 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Nemotron-8B 58.06 26.26 0.00 0.00 0.00 0.09 0.00 0.00 0.00 15.02 0.00 0.57
L-R1-32B 95.07 0.00 0.00 0.00 0.00 0.00 0.81 0.00 0.00 3.03 0.00 1.09
L-R1-32B-DS 99.81 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.19
L-R1-14B-DS 99.19 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.81
L-R1-7B-DS 99.67 0.00 0.05 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
s1.1-32B 99.53 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.43
s1.1-14B 97.39 0.00 0.00 0.14 0.00 0.00 0.24 0.00 0.00 0.00 0.00 2.23
s1.1-7B 88.96 0.00 0.00 7.96 0.09 0.00 0.00 0.00 0.00 0.09 0.00 2.89
EXAONE-32B 67.39 32.42 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.19
EXAONE-7.8B 65.83 32.23 0.00 0.00 0.05 0.47 0.00 0.00 0.00 0.14 0.00 1.28
EXAONE-2.4B 81.42 15.83 0.00 0.09 0.00 0.05 0.00 0.00 0.00 0.05 0.00 2.56
Table 8: Format types rates in top-p sampling setting.
Model Loose Format Error Format
Strict Format
100 101 200 201 202 203 204 205 206 207 300 301
QwQ-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
hunyuan-t1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Sky-T1-32B 99.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.47 0.00 0.24
DS-R1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-70B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-14B 99.76 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
DS-R1-8B 99.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24 0.00 0.24
DS-R1-7B 97.87 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.13
DS-R1-1.5B 91.94 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 8.06
Nemotron-49B 60.90 39.10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Nemotron-8B 55.21 26.78 0.00 0.00 0.00 0.00 0.00 0.00 0.00 16.35 0.00 1.66
L-R1-32B 85.55 0.24 0.00 0.24 0.71 0.24 0.95 0.00 0.00 6.64 2.61 2.84
L-R1-32B-DS 99.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.71
L-R1-14B-DS 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
L-R1-7B-DS 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
s1.1-32B 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
s1.1-14B 95.97 0.00 0.00 0.24 0.00 0.00 0.24 0.00 0.00 0.00 0.00 3.55
s1.1-7B 87.91 0.00 0.00 6.64 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.45
EXAONE-32B 65.88 33.89 0.00 0.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
EXAONE-7.8B 63.51 33.65 0.00 0.00 0.00 0.24 0.00 0.00 0.00 0.24 0.00 2.37
EXAONE-2.4B 78.91 15.88 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.21
Table 9: Format types rates in greedy decoding setting.
C.1 Division of Different Format
This section presents a comprehensive taxonomy of format errors and emphasizes the importance of addressing these issues in future research endeavors. Unlike conventional LLMs, LRMs frequently exhibit format errors, manifested when responses fail to adhere to the requirement of generating a single End Thinking Marker ($\mathtt{ETM}$). This inconsistency creates significant challenges in distinguishing between the thinking process and the final answer. Format errors represent a distinct category from over-thinking and under-accuracy phenomena, yet similarly indicate vulnerability to elementary reasoning traps.
As illustrated in Table 6, we have identified and classified 12 distinct format types, each assigned a unique ID. These format types can be broadly categorized into Strict Format and Loose Format, with Strict Format representing a specialized subset of Loose Format.
Loose Format. ID-100 and ID-101 represent two standard decoding modes, categorized as Strict Format. Both contain only one $\mathtt{ETM}$. Among these, ID-100 includes a thinking process, while ID-101 only generates the $\mathtt{ETM}$ and final answer.
[…]
Error Format. LRMs produce two types of undecodable errors: ID-300 indicates that the LRM outputs only a $\mathtt{ETM}$ without any additional content, while ID-301 represents cases where the LRM’s thinking process reaches the maximum length.
C.2 Evaluation of Different Format
In the classification shown in Table 6, ID-100 / 101 / 201 / 203 / 204 / 206 / 207 have decomposable final answers. If multiple $\mathtt{ETM}$s appear, only the last one is selected for splitting. In this case, the evaluator uses the final answer as the evaluation target.
ID-200 / 202 / 205 lack a final answer, in which case the evaluation target becomes the thinking process, as this part may also contain the correct solution to the question.
C.3 Statistics of Different Format
Table 8 presents the distribution of 12 format types under top-p sampling settings. Our analysis reveals three key insights: (1) The infinite thinking phenomenon is widespread across most models, particularly concentrated in LRMs with fewer than 32B parameters. This indicates that output formatting should be prioritized during training to reduce format errors, especially when working with less powerful foundation models. (2) The Nemotron and EXAONE families frequently produce correctly formatted responses without visible reasoning processes. This behavior can be viewed as a mechanism for mitigating over-thinking. However, the EXAONE family still exhibits substantial overthinking tendencies, suggesting that LRMs’ capability to respond without visible reasoning and their tendency to overthink may be orthogonal characteristics. (3) None of the evaluated LRMs exhibited behaviors classified as ID-205/206, possibly because special $\mathtt{ETM}$s are perceived by LRMs as the only alternative to standard $\mathtt{ETM}$s. Furthermore, ID-300 was not observed, indicating that high-temperature sampling generally yields semantically meaningful responses.
Table 9 shows the format statistics of greedy sampling. The phenomenon is generally similar to top-p sampling. L-R1-32B produced ID-300, possibly due to poor compatibility between the training method and low temperature.
Appendix D Experiment Details
Table 10 shows LRMs’ performance on S1-Bench under greedy sampling settings. While overall accuracy improves compared to top-p sampling settings, $\mathtt{ART}$ slightly decreases, yet overthinking and accuracy degradation issues persist. Figure 9 illustrates the $\mathtt{ART}$ heatmap across 28 subcategories and 26 models.
We compile the statistics of the prejudgment words extracted, as shown in Table 7.
Model Size acc (Loose) acc (Strict) $\mathtt{SUC}$-$\mathtt{L}\uparrow$ $\mathtt{SUC}$-$\mathtt{S}\uparrow$ $\mathtt{ART}$downarrow
QwQ-32B 32B 100.00 100.00 100.00 100.00 750.41
Hunyuan-T1 100.00 100.00 100.00 100.00 541.09
Sky-T1-32B 32B 99.53 99.05 99.76 99.29 157.12
DS-R1 671B 100.00 100.00 100.00 100.00 621.89
DS-R1-70B 70B 99.76 99.76 100.00 100.00 469.78
DS-R1-32B 32B 100.00 100.00 100.00 100.00 428.46
DS-R1-14B 14B 99.29 99.29 99.76 99.76 463.52
DS-R1-8B 8B 97.63 97.39 99.76 99.53 452.11
DS-R1-7B 7B 94.31 94.31 97.87 97.87 436.87
DS-R1-1.5B 1.5B 76.54 76.54 91.94 91.94 473.67
Nemotron-49B 49B 99.53 99.53 100.00 100.00 337.94
Nemotron-8B 8B 84.60 77.73 98.34 81.99 446.62
L-R1-32B 32B 92.18 85.78 94.55 85.78 996.36
L-R1-32B-DS 32B 99.29 99.29 99.29 99.29 528.45
L-R1-14B-DS 14B 98.82 98.82 98.82 98.82 664.28
L-R1-7B-DS 7B 92.65 92.65 98.82 98.82 514.60
s1.1-32B 32B 98.82 98.82 98.82 98.82 983.38
s1.1-14B 14B 95.97 95.50 96.45 95.97 786.30
s1.1-7B 7B 94.31 87.68 94.55 87.91 630.52
EXAONE-32B 32B 97.63 97.39 100.00 99.76 746.89
EXAONE-7.8B 7.8B 86.73 86.49 97.63 97.16 947.92
EXAONE-2.4B 2.4B 72.99 72.99 94.79 94.79 1394.72
Table 10: Main results in the greedy decoding setting on the S1-Bench, sorted by model family. Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst.
Figure 9: $\mathtt{ART}$ on the 28 subcategories, which is the average result of five generations under top-p sampling with Loose Format setting.
Appendix E Prompts discussed in this paper
This section presents the prompts used in this paper. For S1-Bench construction-related prompts, please refer to Table 11; for experimental prompts, please refer to Table 12,13 and 14.
Prompt for construction workflow for S1-Bench
Initial Generation
Based on the category name and explanation, generate 50 pairs of questions and answers in both Chinese and English. Ensure that the generated questions vary in perspective and content, maintaining diversity. All questions must have a **uniquely determined, unambiguous answer with no potential conditions causing multiple possible answers**. Ensure that the questions are expressed clearly. Additionally, the difficulty level of the questions should be **as simple as possible—so easy that they can be answered correctly without any need for thought**.
# Category Name:
{category}
# Explanation:
{explanation}
# Cases:
## English question:
{question_en}
## English Answer:
{answer_en}
## Chinese question:
{question_zh}
## Chinese Answer:
{answer_zh}
Please generate 50 pairs of Chinese and English questions and answers in the following format: [question]English-question[answer]English-answer[question]Chinese-question[answer]Chinese-answer…
Start generating:
Discriminating Generation Quality
Given a question, its answer, and its category, please analyze from the following perspectives as comprehensively as possible:
1. Whether the question belongs to the specified category.
2. Whether the question is clear, unambiguous, has an absolutely unique answer, and has no potential alternative interpretations.
3. Whether the answer is absolutely correct; if not, what the correct answer should be.
4. Whether the question is similar to other given questions, and if similar, whether more diverse questions can be generated.
# Question Category and Explanation:
{category_with_exp}
# Question and Answer:
{question_with_answer}
# Other Questions:
{questions_list}
Begin your analysis, aiming to be as detailed and comprehensive as possible:
Reduce Difficulty Generation
Given a question and answer that are too complex for the model to answer correctly, you need to further reduce their difficulty while trying to:
- Maintain the original perspective and semantic information
- Preserve the question’s alignment with its category
# Question Category and Explanation:
{category_with_exp}
# Question and Answer:
{question_with_answer}
The new question and answer:
Evaluation on S1-Bench
**Question:**
{question}
**Model Answer:**
{model_answer}
**Ground Truth:**
{ground_truth}
Your task is to evaluate whether the model’s answer is correct.
An answer is considered correct as long as it contains the ground truth (regardless of how complex or detailed the description is).
If there are parenthetical notes after the ground truth, then there may be multiple correct answers. In this case, the given answer is just one example, and any answer that meets the requirements specified in the notes can be considered correct.
Additionally, some reasonably uncertain supplementary information is also considered appropriate, including more details, possibilities, and expanded discussion. You should focus more on whether the reply contains the correct answer.
You need to output a standard JSON, providing your explanation of the evaluation in the "explain" field, and giving the evaluation result in the "result" field, where 1 means the answer is correct and 0 means it is incorrect.
Your action should follow the given format: "explain": "", "result": 0/1
Table 11: Prompt for construction workflow for S1-Bench. The “explanation” (and “category_with_exp”) includes the definition of each major category, simplicity requirements in Section 3.1.1, and the definition of subcategories.
Prompts for Solution Segmentation
# Task Description:
Your task is to segment the given Chain of Thought according to the following rules:
1. Segmentation positions:
(1) Please identify and extract all sub-reasoning processes from the Chain of Thought that meet the following condition: They explicitly arrive at a conclusion (including cases phrased as questions, e.g., "right?") that is directly consistent with the Ground Truth. Reasoning processes that only indirectly support the Ground Truth or result in partially aligned conclusions should be excluded.
(2) After clearly reaching the conclusion consistent with the Ground Truth, insert the segmentation marker <split> to distinctly separate each qualifying sub-reasoning process.
2. Output Restriction:
(1) You should only directly output the segmentation result without adding any additional supplements.
(2)Except for inserting the <split> separator, you must not make any other modifications to the original Chain of Thought, not even minor character-level changes such as punctuation, spacing, or capitalization. In other words, after removing all <split> separators, the output text must match the original Chain of Thought exactly.
# Some cases:
## Case 1:
### Question:
What is the value of x if x plus 3 equals 8?
### Ground Truth:
5
### Chain of Thought:
Okay, so I need to figure out what x is if x plus 3 equals 8. The problem says "x plus 3 equals 8," so in mathematical terms, that would be: x + 3 = 8
Alright, now I need to solve for x. From what I remember in algebra, the goal is to isolate x on one side of the equation. That means I need to get rid of that "+ 3" that’s being added to x. To do that, I should perform the opposite operation. Since it’s adding 3, the opposite would be subtracting 3.
So, if I subtract 3 from both sides, that should keep things equal. Let me try that: x + 3 - 3 = 8 - 3
Simplifying both sides, on the left side, the +3 and -3 cancel each other out, leaving just x. And on the right side, 8 minus 3 is 5. So that simplifies to: x = 5
Wait, let me check if that makes sense. If I substitute x back into the original equation, does it hold true? Let’s see:
Original equation: x + 3 = 8
Substitute x = 5: 5 + 3 = 8
5 plus 3 is indeed 8, so that works out.
Another way might be to consider what number, when you add 3 to it, gives 8. So starting from 8, if I take away 3, I should get back to the original number. That’s the same as subtracting 3 from 8, which again gives 5. So that’s consistent with the first method.
Wait a second, what if the original problem was written differently? The problem says "x plus 3 equals 8." In symbols, that’s unambiguous: x + 3 = 8. So no, I think that’s right.
*Final Answer**
The value of x is 555.
### Output:
Okay, so I need to figure out what x is if x plus 3 equals 8. The problem says "x plus 3 equals 8," so in mathematical terms, that would be:x + 3 = 8
Alright, now I need to solve for x. From what I remember in algebra, the goal is to isolate x on one side of the equation. That means I need to get rid of that "+ 3" that’s being added to x. To do that, I should perform the opposite operation. Since it’s adding 3, the opposite would be subtracting 3.
So, if I subtract 3 from both sides, that should keep things equal. Let me try that: x + 3 - 3 = 8 - 3
Simplifying both sides, on the left side, the +3 and -3 cancel each other out, leaving just x. And on the right side, 8 minus 3 is 5. So that simplifies to: x = 5<split>
Wait, let me check if that makes sense. If I substitute x back into the original equation, does it hold true? Let’s see:
Original equation: x + 3 = 8
Substitute x = 5: 5 + 3 = 8
5 plus 3 is indeed 8, so that works out.<split>
Another way might be to consider what number, when you add 3 to it, gives 8. So starting from 8, if I take away 3, I should get back to the original number. That’s the same as subtracting 3 from 8, which again gives 5.<split> So that’s consistent with the first method.
Wait a second, what if the original problem was written differently? The problem says "x plus 3 equals 8." In symbols, that’s unambiguous: x + 3 = 8. So no, I think that’s right.
*Final Answer**
The value of x is 555.<split>
## Case 2:
## Case 3:
# Eval Target:
## Question:
{question}
## Ground Truth:
{ground_truth}
## Chain of Thought:
{thinking_process}
## Output:
Table 12: Prompts for Solution Segmentation.
Prompts for Error Analysis
Types of Correct
# Task Description:
You will receive a Question, its corresponding Ground Truth, and a Chain of Thought(COT) generated by a LLM for that Question.
Your task is to carefully analyze the CoT and assign it to one of the two predefined categories listed below.
# Categories:
1: The CoT ***includes explicit incorrect conclusions*** in intermediate reasoning steps.
2: The CoT ***doesn’t include any explicit incorrect conclusion*** in intermediate reasoning steps.
Output your evaluation in the following format:
# TheReason: [note: Conduct a step-by-step analysis, stating if and where explicit incorrect conclusions occur in the COT.]
# ErrorType: [note: Summarize each incorrect conclusion into a specific error type using a phrase of less than 5 words, such as factual inaccuracies, logical fallacies, comprehension mistakes, calculation errors, formatting issues, and so forth, to better conduct further evaluation and analysis. Directly output a Python list, where each element represents the error type of a specific incorrect conclusion in the CoT. If there are no incorrect conclusions, return an empty list.]
# TheCategory: [note: Provide your classification based on your analysis using only the number "1" or "2". Do not add any additional text.]
# Question:
{question}
# Ground Truth:
{gound_truth}
# COT:
{thinking_process}
# TheReason:
Types of Error
# Task Description:
You will receive a Question, its corresponding Ground Truth, and a Chain of Thought(COT) generated by a LLM for that Question.
Your task is to carefully analyze the CoT and assign it to one of the two predefined categories listed below.
# Categories:
1: Regardless of whether the CoT ultimately arrives at the correct final answer or not, ***the correct answer is explicitly mentioned at least once*** within the reasoning steps (even if it is not ultimately adopted).
2: ***The correct answer is never explicitly mentioned or referenced*** at any point within the reasoning steps.
Output your evaluation in the following format:
# TheReason: [note: Conduct a step-by-step analysis, explicitly stating whether and where a correct answer is mentioned within the reasoning steps.]
# TheCategory: [note: Provide your classification based on your analysis using only the number "1" or "2". Do not add any additional text.]
# Question:
{question}
# Ground Truth:
{answer}
# COT:
{thinking_part}
# TheReason:
Table 13: Prompts for Error Analysis.
Prompts for Extracting Prejudgment Sentences and Words
Extracting Sentences
# Given the beginning of a model’s thought process when answering a question, determine if there is any evaluation by the model regarding the difficulty level or category of the question.
# For example:
Example Evaluation 1: This seems to be an easy question.
Example Evaluation 2: {CJK}UTF8gkai这看起来似乎是个基础的问题
Example Evaluation 3: This question is a bit complex.
Example Evaluation 4: {CJK}UTF8gkai这个问题看起来复杂
# Beginning of thought process:
{thinking_process}
# Please output only the evaluative short phrase you find. If there is no evaluation, output "None".
# Output:
Extracting Words
# You are given a comment sentence, and your goal is to find the only comment-related word.
# For example:
## case 1:
Sentence: {CJK}UTF8gkai这是一个很简单的算术题
Word: {CJK}UTF8gkai简单
## case 2:
Sentence: that seems pretty straightforward
Word: straightforward
# The comment sentence:
{review_sentence}
Output the only comment-related word:
Table 14: Prompts for Extracting Prejudgment Sentences and Words.
cate. subcatrgories Explanation and cases
reasoning question numerical reasoning Questions that require performing basic mathematical operations or solving simple algebraic
equations to arrive at a numerical answer.
Case: What’s two plus three?
code reasoning Questions that require tracing through and executing simple code snippets to determine their
output or behavior when run in a specific programming environment.
Case: What is the output of the following code when run in Python 3 environment: word =
[…]
set reasoning Questions that require applying simple syllogistic reasoning to determine whether elements
belong to sets based on clearly stated relationships.
Case: All squares are quadrilaterals. A shape is a square, is it a quadrilateral?
temporal reasoning Questions that require calculating time durations, ages, or future dates by applying simple
arithmetic operations to temporal information.
Case: How many minutes equal 120 seconds?
spatial reasoning Questions that require determining relative positions, directions, or orientations of objects
in space based on simple spatial relationships.
Case: If a bird is flying above a tree, where is the tree in relation to the bird?
causal reasoning Questions that require determining outcomes by applying simple cause-and-effect
relationships based on given conditional statements.
Case: If ferromagnetic material is placed in a magnetic field, it will become magnetized. An
iron nail was placed next to a strong magnet for some time. Has the nail been magnetized?
natural law reasoning Questions that require applying basic knowledge of physical laws and natural phenomena to
predict simple observable outcomes in everyday scenarios.
Case: Which is faster, an airplane or the propagation of light?
knowledge question geometry facts Questions that require recalling simple and fundamental geometric properties about shapes,
angles, and basic geometric figures.
Case: How many angles does a trapezoid have?
geographic facts Questions that require recalling simple factual information about locations, landmarks,
political divisions, celestial bodies, and other basic geographic knowledge.
Case: Which is the largest continent on Earth?
historical facts Questions that require recalling basic facts about historical events.
Case: Which country first invented paper?
biographical facts Questions that require recalling basic facts about the identities, achievements, and
characteristics of historical figures.
Case: Who proposed the theory of universal gravitation?
measurement units Questions that require recalling simple conversion relationships between
standard units of measurement.
Case: How many centimeters equal 1 meter?
scientific notation Questions that require recalling basic scientific symbols, formulas, and standard units
used in scientific communication.
Case: What is the chemical symbol for oxygen?
creative authorship Questions that require recalling the creators or originators of notable artistic, literary,
musical, and cultural works.
Case: Who is the author of Hamlet?
Table 15: The subcategory descriptions and cases of reasoning questions and knowledge questions.
cate. subcatrgories Explanation and cases
instruction following repetition constraints Questions that require outputting specified characters, words, or phrases a specific number
of times according to simple formatting instructions.
Case: Output the number "7" four times, without using separators.
length constraints Questions that require generating outputs of a specific length or with a specific number
of components based on simple counting constraints.
Case: Output a four-digit number.
character constraints Questions that require generating words or numbers that conform to simple specified
character patterns or formatting rules.
Case: Output a number that begins with 8.
counting constraints Questions that require counting specific characters or elements within
a given text or sequence.
Case: Output the number of letter "y" in the word "yes".
transformation constraints Questions that require modifying text or numbers according to simple formatting or character
substitution rules to produce a transformed output.
Case: Output the word "good" with all letters capitalized directly.
sentence constraints Questions that require generating sentences that conform to simple specified content or
structural requirements.
Case: Give a sentence that contains the phrase "have lunch" directly.
analysis question sentiment classification Questions that require determining whether simple statements express positive or negative
emotions based on the tone and word choice.
Case: Does the sentence "I hate rainy days." express a positive or negative emotion?
named entity recognition Questions that require identifying the correct category of named entities (such as people,
places, organizations, or time expressions) within simple sentences.
Case: In the sentence "Napoleon died in 1821", is "1821" a time or a place name?
language classification Questions that require identifying the language of origin for simple words or phrases
based on their characteristic writing systems or common vocabulary.
Case: Is the word "hello" English or Japanese?
topic classification Questions that require identifying the primary subject matter or thematic category of
simple sentences based on their content and context clues.
Case: Is the topic of the sentence "The stock market rose 2% today" finance or technology?
intent recognition Questions that require determining the communicative purpose behind simple utterances
or statements based on their phrasing and context.
Case: Is the intention of the sentence "I’m sorry I’m late." to apologize or to blame?
syntax classification Questions that require identifying the correct grammatical structure or sentence type of
simple expressions based on their form, punctuation, and communicative function.
Case: Is "Close the door!" an imperative sentence or an interrogative sentence?
grammar classification Questions that require identifying simple grammatical properties (like tense, voice,
or polarity) of sentences based on their structure and verb forms.
Case: Is "The apple was eaten." in active voice or passive voice?
coreference resolution Questions that require identifying which entity a pronoun or reference term refers to in
simple sentences by tracking relationships between words in the text.
Case: In "My computer is broken, and I need to fix it." What does "it" refer to?
Table 16: The subcategory descriptions and cases of instruction following questions and analysis questions.

### v3

@@
 Abstract
-We introduce S1-Bench, a novel benchmark designed to evaluate Large Reasoning Models’ (LRMs) performance on simple tasks that favor intuitive system 1 thinking rather than deliberative system 2 reasoning.
-While LRMs have achieved significant breakthroughs in complex reasoning tasks through explicit chains of thought, their reliance on deep analytical thinking may limit their system 1 thinking capabilities.
-Moreover, a lack of benchmark currently exists to evaluate LRMs’ performance in tasks that require such capabilities.
-To fill this gap, S1-Bench presents a set of simple, diverse, and naturally clear questions across multiple domains and languages, specifically designed to assess LRMs’ performance in such tasks.
-Our comprehensive evaluation of 22 LRMs reveals significant lower efficiency tendencies, with outputs averaging 15.5 times longer than those of traditional small LLMs.
-Additionally, LRMs often identify correct answers early but continue unnecessary deliberation, with some models even producing numerous errors.
-These findings highlight the rigid reasoning patterns of current LRMs and underscore the substantial development needed to achieve balanced dual-system thinking capabilities that can adapt appropriately to task complexity [^1] The code and benchmark can be found in https://github.com/WYRipple/S1_Bench..
+We introduce S1-Bench, a novel benchmark designed to evaluate the performance of Large Reasoning Models (LRMs) on simple tasks that favor intuitive system 1 thinking rather than deliberative system 2 reasoning.
+While LRMs have achieved significant breakthroughs in complex reasoning tasks through explicit chains of thought, their heavy reliance on system 2 thinking may limit their system 1 thinking capabilities.
+However, there is a lack of an appropriate benchmark for evaluating LRM’s system 1 thinking capabilities.
+To fill this gap, S1-Bench introduces a suite of simple, diverse, and natural questions across multiple domains and languages, specifically designed to assess LRMs’ performance on questions more suitable for system 1.
+We conduct extensive evaluations across 28 LRMs, revealing their inefficiency, inadequate accuracy, and limited robustness when handling simple questions.
+Additionally, we observe a gap between their difficulty perception and generation length.
+Overall, this work paves the way toward dual-system compatibility in the development of LRMs [^1] The code and benchmark can be found in https://github.com/WYRipple/S1_Bench..
 […]
@@
 denotes corresponding author., Shuaiyi Nie, Xinghua Zhang, Zefeng Zhang, Tingwen Liu Institute of Information Engineering, Chinese Academy of Sciences School of Cyber Security, University of Chinese Academy of Sciences {zhangwenyuan,nieshuaiyi,liutingwen}@iie.ac.cn
-Figure 1: LRMs exhibit under-accuracy and overthinking on simple problems.
-Shapes represent organizations, colors represent base model families, with darker colors indicating larger models, and connecting lines represent the relationships between model families and training.
 1 Introduction
@@
 Recent advances in Large Reasoning Models (LRMs), notably OpenAI’s o1/o3 OpenAI (2024) and the DeepSeek-R1 Guo et al.
-(2025) series, have propelled the development of Large Language Models (LLMs) by explicitly generating external chain-of-thought (COT) Wei et al.
-(2022) before producing final answers, achieving remarkable performance on complex reasoning tasks.
-Unlike traditional LLMs that primarily rely on intuitive, heuristic system 1 thinking, LRMs enhance their performance by employing analytical, deliberative system 2 thinking Qu et al.
-(2025); Li et al.
-(2025b), which includes sophisticated strategies such as self-reflection, multi-path exploration Li et al.
+(2025) series, have propelled the development of Large Language Models (LLMs).
+Unlike traditional LLMs that exhibit intuitive, heuristic system 1 thinking, LRMs demonstrate deliberate and analytical system 2 reasoning Qu et al.
+(2025a); Li et al.
+(2025b) by explicitly generating external chain-of-thought (COT) Wei et al.
+(2022) before producing final answers.
+Through sophisticated strategies such as self-reflection and multi-path exploration Li et al.
 (2025a); Yeo et al.
+(2025), LRMs can achieve strong performance in tasks that require system 2 thinking, including advanced mathematical and competition-level problems Yang et al.
+(2025a).
+Benchmark Cross Realistic Multi- Acc.
+Domain Scenarios lingual
+AIME 6.67
+GPQA 24.94
+Olympiad-Bench 27.94
+AMC 31.88
+MATH 58.30
+MMLU 66.27
+GSM8K 87.45
+ASDIV 97.51
+GSM8K-zero 77.98
+RoR-Bench 14.24
+S1-Bench (ours) 100.00
+Table 1: Characteristics of S1-Bench, with “Acc.” representing the average accuracy of four 7-9B LLMs. See Appendix A.1 for more details.
+However, there remains a lack of appropriate benchmarks for evaluating LRMs’ system 1 thinking capabilities.
+Not all real-world problems require system 2 reasoning.
+The capacity to dynamically identify simple questions and address them efficiently contributes to both resource optimization and improved user satisfaction for LRMs.
+Nevertheless, current benchmarks either overemphasize difficulty, are simple yet lack domain diversity, or are only not hard for humans but involve unrealistic adversarial designs.
+Table 1 presents a collection of recent benchmarks aimed at mitigating the overthinking in LRMs Sui et al.
 (2025).
-Despite the benefits of system 2 thinking, a rigid reliance on this cautious thinking mode can also introduce new limitations, such as performance degradation in specific contexts Feng et al.
-(2025); Zhao et al.
-(2025) or overthinking Chen et al.
-(2024); Kumar et al.
-(2025a).
-However, a research gap exists in thoroughly investigating how LRMs’ over-reliance on system 2 thinking affects their performance when confronting extremely simple questions better suited for intuition-driven system 1 processing.
-Specifically, existing research primarily analyzes LRM behaviors in complex tasks that naturally align with deliberative system 2 thinking, such as MATH for math problems Luo et al.
-(2025); Yang et al.
-(2025); Chen et al.
-(2024) or IFEval for instruction following Zhao et al.
-(2025).
-Although some benchmarks involve relatively simple questions for human Miao et al.
-(2021); Yan et al.
-(2025); Chiang and Lee (2024), these tasks are often adversarial or non-trivial for LLMs, requiring moderate reasoning and can not be effectively solved through intuitive system 1 processing.
+The majority of these benchmarks are of high difficulty.
+For example, AIME and GPQA MAA Committees ; Rein et al.
+(2024) achieve less than 30% accuracy on conventional small LLMs, which are inherently more suitable for system 2 reasoning.
+Although some simple mathematical benchmarks are easy enough, such as GSM8K and ASDIV Cobbe et al.
+(2021); Miao et al.
+(2021), they often suffer from limited domain variety.
+Furthermore, some tasks that pose little challenge to humans but incorporate adversarial elements tend to lack relevance to realistic scenarios, such as GSM8K-zero Chiang and Lee (2024), which includes the correct answer in the questions.
 Thus, a benchmark to assess the system 1 thinking capability of LRMs is still lacking, further hindering our understanding of LRMs’ cognitive flexibility between the two systems Ziabari et al.
 (2025); Qu et al.
-(2025).
+(2025a).
 To fill this research gap, we introduce the System 1 Thinking Capability Benchmark (S1-Bench), which measures the performance of LRMs across various simple tasks that commonly encountered in real-world applications.
@@
 (2) Diverse.
-S1-Bench is not limited to simple reasoning problems; it encompasses four major categories and 28 subcategories in two languages (English and Chinese), including commonsense knowledge, loosely constrained instruction following, and straightforward analytical problems.
+S1-Bench is not limited to simple reasoning problems; it encompasses four major categories and 28 subcategories in two languages (English and Chinese), including reasoning problems, commonsense knowledge, instruction following, and analytical problems.
 (3) Natural.
 The questions are clear, without any misleading elements or ambiguities, ensuring they can be answered intuitively.
-S1-Bench conducted extensive evaluations across 22 LRMs, yielding the following key findings: (1) Current LRMs exhibit inefficiency and lack system 1 thinking capabilities, with average output lengths 15.5 times longer than small LLMs.
-(2) LRMs often arrive at correct answers early in their thinking process but continue unnecessary reasoning with gradually increasing content similarity.
-(3) Despite employing deep reasoning, LRMs exhibit accuracy degradation on simple questions compared to traditional LLMs.
-(4) LRMs show the ability to prejudge the simplicity of questions, yet they still exhibit inefficiency.
+We conduct extensive evaluations on S1-Bench across 28 LRMs, yielding the following key findings: (1) Current LRMs exhibit inefficiency and lack system 1 thinking capabilities across all types of questions, with average output lengths 15.5 times longer than small LLMs on S1-Bench.
+(2) Despite employing deep reasoning, several LRMs exhibit under-accuracy and limited robustness on simple questions.
+(3) LRMs exhibit “gut moment” at the beginning of some reasoning processes, showing gut feelings about task difficulty.
+Yet, even when recognizing a question’s simplicity, LRMs often fail to produce shorter responses—revealing a gap between their difficulty awareness and generation behavior.
 These findings emphasize the significant distance LRMs must traverse to become powerful dual-system compatible models.
 Our contributions can be summarized as follows:
-  • To the best of our knowledge, S1-Bench is the first benchmark to evaluate the system 1 thinking capabilities of LRMs.
-  • We introduce a workflow for constructing simple dataset suitable for system 1 evaluation.
-  • Extensive experiments reveal the inefficiency and under-accuracy of LRMs on simple questions that are suited for system 1 thinking.
+  • To the best of our knowledge, S1-Bench is the first benchmark to evaluate the system 1 thinking capabilities of LRMs, which paves the way for dual-system compatibility.
+  • We introduce a workflow for constructing a simple dataset for system 1 evaluation.
+  • Extensive experiments reveal the inefficiency, under-accuracy, and limited robustness of LRMs on simple questions.
+  • We find that LRMs exhibit “gut moment” on simple problems, and reveal a gap between their difficulty perception and generation length.
 2 Related work
@@
 (2025b); Chen et al.
-(2025), achieve a paradigm shift from intuitive system 1 thinking to deliberative system 2 reasoning compared to traditional LLMs Li et al.
+(2025a), achieve a paradigm shift from intuitive system 1 thinking to deliberative system 2 reasoning compared to traditional LLMs Li et al.
 (2025b); Qu et al.
-(2025), thus achieving superior performance on complex tasks.
+(2025a), thus achieving superior performance on complex tasks.
 The development of recent LRMs has largely followed two main approaches: large-scale reinforcement learning (RL) and model distillation.
@@
 (2025); Team (2025b); Team et al.
-(2025) leverage reward-based optimization to gradually incentivize deliberative reasoning.
+(2025b) leverage reward-based optimization to gradually incentivize deliberative reasoning.
 In contrast, distillation-based LRMs OpenAI (2024); Min et al.
@@
 (2025); Muennighoff et al.
-(2025) acquire such abilities by transferring structured reasoning patterns from advanced teacher models.
+(2025); Zhang et al.
+(2025b) acquire such abilities by transferring structured reasoning patterns from advanced teacher models.
 2.2 Limitations of LRMs
@@
 On one hand, intermediate reasoning steps can cause excessive token generation and unnecessary solving attempts Chen et al.
-(2024); Hashemi et al.
+(2024b); Hashemi et al.
 (2025); Kumar et al.
-(2025a).
+(2025a), even leading to redundancy in the hidden layers Chen et al.
+(2024c, 2025b).
 On the other hand, LRMs’ performance can drop in specific contexts like safety scenarios Jiang et al.
-(2025); Zhao et al.
 (2025) and role-playing Feng et al.
@@
 3.1 How to Ensure Simplicity?
-We ensure questions are simple and suitable for system 1 thinking through two aspects.
+We ensure questions are simple and suitable for system 1 thinking through the following two aspects.
 3.1.1 A Priori Simplicity Constraints
-We require all questions and answers to adhere to the following guidelines:
-(1) Questions must be naturally, clearly expressed, and unambiguous, free of intentional traps.
-(2) Answers must be unique or easily falsifiable (such as providing a three-letter English word).
-We’ve further implemented category-specific constraints to ensure simplicity: Rsn: Limited to problems solvable with minimal reasoning or intuition.
+We begin by generating question–answer pairs through collaboration between humans and LLMs. Each pair is required to satisfy both the general and the category-specific simplicity criteria.
+The general simplicity criteria requires that: (1) Questions must be naturally and clearly expressed, unambiguous, and free of intentional traps.
+(2) Answers must be unique or easily falsifiable (e.g., providing a three-letter English word).
+The category-specific simplicity criteria are as follows.
+Rsn: Limited to problems solvable with minimal reasoning or intuition.
 Kno: Restricted to common knowledge with unique, verifiable answers from sources like Wikipedia.
@@
 Therefore, we introduce additional posteriori verification to ensure that questions are simple enough to be correctly and robustly answered by smaller LLMs from different families.
-This alleviates cognitive biases in traditional LLMs for system 1 thinking problems Li et al.
-(2025b).
 3.2 Construction Workflow
-To implement a priori constraints, we introduce data generators [^2] We select Claude-3.7-Sonnet and Qwen2.5-72B-Instruct.
-and quality discriminators [^3] We select GPT-4o and DeepSeek-V3-241226.
-that adhere to the a priori settings.
-To ensure posterior verification, we introduce simplicity validators [^4] We select four small LLMs: Qwen2.5-7B, Llama3.1-8B, Mistral-8B, and Gemma2-9B.
-The full model IDs are detailed in Table 4.
-and correctness evaluator [^5] We select GPT-4o..
-The generators, discriminators, and evaluator are uniformly configured with temperature 0.7 and top-p=0.9, while validators are set with temperatures of 0.0, 0.2, and 0.4, each sampling 10 times.
-Additionally, we recruit three experienced graduate students as annotators, familiar with LLMs and fully comprehending the construction objectives of S1-Bench.
-For the prompts used in the workflow, refer to Table 11.
 Subcategory Preparation.
-Initially, to ensure diversity, we draw on existing benchmark classification methods (e.g., MMLU, IFEval and GSM8K) and evaluation surveys Chang et al.
-(2024), selecting or merging existing subcategories, and designing new ones.
-We ensure that all subcategories meet the requirements for simplicity and provide an example question for each.
-Generation and Modification.
-We use two data generators to create initial questions and answers for each candidate subcategory.
-The generation process adheres to prompt requirements (refer to Table 11), ensuring bilingual content that aligns with subcategory definitions while maintaining diverse perspectives.
-Each question is accompanied by auxiliary answers, with 50 initial bilingual question-answer pairs generated per subcategory.
-These pairs are independently assessed by three annotators and two quality discriminators , who ensure that the questions are unambiguous, have clear boundaries, and admit only one correct answer.
-For each question-answer pair, five independent assessments are conducted.
-Based on these evaluations, annotators discuss and collectively decide whether to retain, modify, or discard each question, ensuring alignment with the requirements outlined in Section 3.1.1.
-Retained questions are submitted to validators for testing, and an evaluator assesses the accuracy of responses.
-Questions with correct answers are included in S1-Bench, while others undergo reprocessing.
-Questions with errors in any of the 30 sampling iterations undergo an iterative difficulty reduction process.
-These questions return to generators with specific prompts to decrease complexity, followed by the same discrimination, discussion, and evaluation procedures.
-Questions failing to achieve error-free performance after three difficulty reduction iterations are excluded from the workflow.
-The final S1-Bench comprises questions that satisfy both human-established a priori simplicity constraints and a posteriori simplicity verification.
+To ensure diversity, we refer to the subcategories included in existing benchmarks (e.g., MMLU, IFEval, and GSM8K) and evaluation surveys Chang et al.
+(2024) to select, merge, or design subcategories for S1-bench, ensuring that each meets the simplicity requirements.
+The definition and example question for each subcategory can be found in Appendix A.2.
+Implementation of A Priori Simplicity.
+First, we use two data generators [^2] We select Claude-3.7-Sonnet and Qwen2.5-72B-Instruct.
+to create 100 initial bilingual question-answer pairs for each candidate subcategory.
+The data generation prompt (see Appdiix A.3) explicitly incorporates the subcategory definitions, along with both the general and category-specific simplicity criteria, while also aiming to ensure diversity in the generated questions.
+Second, these question–answer pairs are then independently evaluated by three annotators and two quality discriminators [^3] We select GPT-4o and DeepSeek-V3-241226.
+according to the general and category-specific simplicity criteria (see Appdiix A.3 for prompt of discriminators), resulting in five evaluation outcomes per pair.
+The three annotators are experienced graduate students familiar with LLMs and well-acquainted with the goals of S1-Bench.
+Finally, based on these evaluation outcomes, three annotators discuss and collectively decide whether to retain, modify, or discard each question.
+Implementation of A Posteriori Simplicity.
+First, each question obtained from the previous stage is input into the small LLM validators [^4] We select four small LLMs: Qwen2.5-7B, Llama3.1-8B, Mistral-8B, and Gemma2-9B.
+The full model IDs are detailed in Table B.1.
+with 7~9 B parameters.
+For each question, we sample 10 answers at three different temperature settings (0, 0.2, and 0.4), resulting in a total of 30 responses per question.
+These responses are then individually evaluated for correctness using GPT-4o.
+Second, if all 30 sampled responses are correct, the question is accepted into S1-Bench.
+Otherwise, the question is returned to the generators, where a difficulty-reduction prompt (see Appendix 10) is applied to simplify it.
+The simplified questions then undergoes the same subsequent process.
+Finally, questions fail to meet the full-accuracy criterion (i.e., 30 out of 30 correct) after three rounds of difficulty reduction are excluded from the workflow.
+The final S1-Bench comprises questions that satisfy both human-based a priori simplicity constraints and LLM-based a posteriori simplicity verification.
 Figure 2: Statistical distribution of token counts for S1-Bench questions.
@@
 Qwen2.5-7B 100.00 100.00 100.00 42.81
+DeepSeek-v3 100.00 100.00 100.00 79.53
+Llama3.3-70B 100.00 99.76 99.76 53.71
 Qwen2.5-14B 99.74 99.76 99.76 40.00
@@
 Qwen2.5-72B 100.00 100.00 100.00 44.61
-Llama3.3-70B 100.00 99.76 99.76 53.71
-DeepSeek-v3 100.00 100.00 100.00 79.53
-Table 1: Average accuracy and response token count of different LLMs, each sampled 10 times at three temperature settings.
+Qwen3-32B (w/o think) 100.00 100.00 100.00 103.30
+Qwen3-14B (w/o think) 100.00 100.00 100.00 86.35
+Qwen3-8B (w/o think) 100.00 100.00 99.76 90.54
+Qwen3-1.7B (w/o think) 98.10 97.16 95.73 114.32
+Table 2: Average accuracy (acc@k) and response token count of different LLMs, each sampled 10 times at three temperature settings.
 3.3 Benchmark Statistics
 S1-Bench comprises 422 question-answer pairs across four major categories and 28 subcategories, balanced with 220 English and 202 Chinese questions.
-We present the category distribution and descriptions with examples for all subcategories, which can be found in Figure 8 and Table 15,16 in the appendix.
 Figure 2 shows the token length distribution, with questions averaging 14.46 tokens.
-To validate the absence of response bias from validators, we tested S1-Bench on five LLMs that serve as base models for commonly used LRMs.
-As shown in Table 1, these models achieved an accuracy exceeding 99% under three temperature settings, confirming the simplicity and robustness of S1-Bench.
-Model ID Size Loose Format Strict Format L-Corr $\uparrow$ S-Corr $\uparrow$ $\mathtt{ART}\downarrow$
+To ensure that the a posteriori verification process does not introduce simplicity only tailored to the small LLM validator, we evaluate S1-Bench on five additional LLMs and on Qwen3 Family with reasoning modes disabled.
+As shown, even the 1.7B model achieves over 98% accuracy at temperature 0.
+Model ID Size Loose Format Strict Format L-Corr $\uparrow$ S-Corr $\uparrow$ Tokens $\downarrow$
 pass@1$\uparrow$ acc@$\uparrow$ pass@1$\uparrow$ acc@$\uparrow$
 Validator LLMs 7-9B 100.00 100.00 100.00 100.00 42.00
+Qwen3-A22B 235B 99.91 99.76 99.91 99.76 100.00 100.00 701.65
+Qwen3-A3B 30B 99.95 99.76 99.95 99.76 100.00 100.00 638.40
 QwQ-32B 32B 100.00 100.00 100.00 100.00 100.00 100.00 720.10
+Qwen3-32B 32B 99.91 99.53 99.91 99.53 99.91 99.91 668.69
+Qwen3-14B 14B 99.95 99.76 99.95 99.76 99.95 99.95 582.99
+Qwen3-8B 8B 99.95 99.76 99.95 99.76 99.95 99.95 657.76
+Qwen3-1.7B 1.7B 99.34 97.39 99.34 97.39 99.81 99.81 595.90
 Hunyuan-T1 99.91 99.53 99.91 99.53 100.00 100.00 542.31
-Sky-T1-32B 32B 98.82 94.79 94.88 79.62 99.48 95.26 163.00
 DS-R1 671B 100.00 100.00 100.00 100.00 100.00 100.00 646.40
@@
 DS-R1-1.5B 1.5B 81.47 54.50 81.47 54.50 97.58 97.58 489.54
+Sky-T1-32B 32B 98.82 94.79 94.88 79.62 99.48 95.26 163.00
 Nemotron-49B 49B 99.15 97.39 99.15 97.39 100.00 100.00 362.54
@@
 4.1 Baseline Models and Configurations
-We evaluated 22 different LRMs, which were explicitly trained to first respond with a thinking process, and then generate a final answer.
+We evaluated 28 different LRMs, which are explicitly trained to first respond with a thinking process, and then generate a final answer.
 These LRMs include open-source families, such as DeepSeek Guo et al.
-(2025), Qwen Team (2025b), Nemotron Chris Alexiuk and Patel (2025), Light-R1 Wen et al.
+(2025), Qwen Yang et al.
+(2025a), Nemotron Bercovich et al.
+(2025), Light-R1 Wen et al.
 (2025), s1.1 Muennighoff et al.
@@
 (2025), and Sky-T1 Griggs et al.
-(2025), as well as closed-source Hunyuan-T1 Tencent (2025), spanning from tiny (1.5B) to large (671B) parameter sizes [^6] More details in Appendix Table 5 (includes training data scale and algorithms)..
+(2025), as well as closed-source Hunyuan-T1 Tencent (2025), spanning from tiny (1.5B) to large (671B) parameter sizes [^5] Model details are presented in Appendix B.1..
 Notably, OpenAI’s o-series models are not included as they do not disclose thinking processes to users.
 For each model, we consider two sets of generation configurations: Greedy sampling with temperature t=0; Top-p sampling with temperature t=0.6, top-p=0.95 and sampling size k=5.
+Only top-p sampling results are reported in the main text; greedy decoding results are provided in the Appendix C.1.
 4.2 Evaluation Metrics
 Format Metrics.
-Generally, the output of LRMs should be separated by an “end thinking marker” (e.g., </think>) that distinguishes the thinking process from the final answer.
-However, LRMs do not always generate responses in correct format.
-To assess this, we calculate the percentage of responses that meet the required formatting criteria, categorized as follows (for top-p sampling setting, each metric is averaged across 5 generations):
-S-Corr (Strict Format Correctness Rate): The percentage of responses that strictly adhere to the format, meaning they only contain one "end thinking marker" followed by a non-empty final answer.
-L-Corr (Loose Format Correctness Rate): The percentage of responses that loosely adhere to the format, meaning they meet the format requirements except for responses with endless repetitions [^7] Detailed format analyses are in the Appendix C..
+To assess the formatting quality of LRM responses, we compute the proportion of responses that satisfy the following two formatting criteria (averaged over 5 runs for top-p sampling).
+S-Corr (Strict Format Correctness Rate): In general, an end thinking marker (e.g., </think>) is expected to separate the thinking process from the non-empty final answer.
+S-Corr measures the proportion of responses that satisfy this criterion.
+L-Corr (Loose Format Correctness Rate): LRMs may occasionally generate responses with endless thinking. L-Corr quantifies the proportion of responses that do not exhibit this failure mode.
+Detailed format types are given in Appendix B.4.
 Model ID Size S1-Bench-$\mathtt{EN}$ S1-Bench-$\mathtt{ZH}$ Avg
@@
 Hunyuan-T1 561.6 693.8 380.9 435.0 521.2 676.8 553.8 505.1 523.8 565.3 542.3
+Qwen3-14B 14B 700.4 639.5 286.2 575.0 579.8 730.4 557.2 403.1 586.0 586.5 583.0
+Qwen3-1.7B 1.7B 790.4 720.6 399.9 526.2 624.6 689.8 563.6 406.4 545.9 564.7 595.9
+Qwen3-A3B 30B 745.0 729.3 328.1 594.8 625.7 773.7 655.8 453.7 648.6 652.2 638.4
 DS-R1 671B 786.1 723.8 711.4 529.2 672.5 727.3 638.5 607.9 533.9 617.9 646.4
+Qwen3-8B 8B 853.7 753.1 394.4 629.5 683.2 749.2 623.8 459.3 624.0 630.0 657.8
+Qwen3-32B 32B 805.7 774.2 356.9 645.5 674.7 780.2 695.2 446.6 645.3 662.1 668.7
 L-R1-14B-DS 14B 951.0 1026.0 829.8 653.5 848.2 594.7 610.1 442.2 451.7 525.7 693.2
+Qwen3-A22B 235B 925.3 864.3 487.2 605.7 734.5 803.3 713.4 487.2 611.3 665.9 701.7
 s1.1-7B 7B 1039.5 840.8 1923.2 529.4 929.9 489.6 351.3 1034.3 332.4 475.6 711.5
@@
 EXAONE-2.4B 2.4B 1927.3 1426.2 1200.1 825.7 1320.7 2469.7 1622.6 2471.6 1511.2 1898.7 1594.0
-Column Avg - 810.6 771.2 896.8 590.2 736.0 679.8 521.6 742.0 447.1 563.0 652.8
-Improvement - $\times$10.9 $\times$18.5 $\times$109.7 $\times$11.0 $\times$14.9 $\times$14.1 $\times$13.8 $\times$62.7 $\times$14.8 $\times$16.7 $\times$15.5
-Table 3: $\mathtt{ART}$ on the S1-bench across two languages and four main categories.
+Column Avg - 809.1 766.0 785.1 591.5 718.4 695.8 545.9 677.9 482.0 576.7 650.3
+Improvement - $\times$10.8 $\times$18.4 $\times$96.0 $\times$11.1 $\times$14.5 $\times$14.4 $\times$14.5 $\times$57.3 $\times$16.0 $\times$17.1 $\times$15.5
+Table 4: Average response tokens in the top-p sampling setting on the S1-bench across two languages and four main categories.
 Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst.
@@
 Efficiency Metrics.
-We calculate the Average Response tokens (ART), which represent the average number of tokens in responses that meet the loose formatting requirements.
-Token counts are obtained using the Qwen2.5 tokenizer.
+We calculate the average token counts for responses (Tokens) except for those generate endless thinking. Token counts are obtained using the Qwen2.5 tokenizer.
 Accuracy Metrics.
-We calculate accuracy metrics for both strict and loose formatting requirements, respectively.
-We use GPT-4o as the evaluator to assess the correctness of the responses [^8] If a final answer can be isolated, only the final answer is evaluated; otherwise, the entire response is assessed., with the evaluation prompt in Table 11.
-For greedy sampling, we directly calculate the accuracy rate, with complete results in Appendix D.
-For top-p sampling, we utilize two metrics:
+We calculate accuracy metrics under both strict and loose formatting requirements, respectively.
+We use GPT-4o as the evaluator to assess the correctness of the responses [^6] If a final answer can be isolated, only the final answer is evaluated; otherwise, the entire response is assessed., with the evaluation prompt in Appendix B.2.
+For greedy sampling, we directly calculate the accuracy rate.
+For top-p sampling, we utilize two metrics: Pass@1 and Acc@k.
+Pass@1 follows DeepSeek-R1 Guo et al.
+(2025), and Acc@k is the percentage of questions with all k answers correct.
+The two metrics use k=5, and their detailed definitions can be found in Appendix B.3.
+Notably, S-Corr \ L-Corr represents the upper bound for pass@1 and acc@5 under strict \ loose formatting requirements.
+4.3 Main Results
+Table 3 and Figure 8 presents the main results of LRMs on S1-Bench, revealing two key phenomena.
+LRMs exhibit significantly lower efficiency than LLMs on S1-Bench, and no clear correlation is observed between ART and model size.
+We observed the following: First, state-of-the-art LRMs, such as DeepSeek-R1 and Qwen3, do not demonstrate a distinct advantage in efficiency.
+In contrast, Sky-T1-32B, which undergoes specific optimizations to mitigate overthinking using SimPO, achieves the highest efficiency.
+Second, The L-R1-DS 7B/14B/32B models are further post-trained from the DS-R1-7B/14B/32B models.
+The L-R1-DS models tend to produce longer responses, suggesting that while additional post-training may enhance the model’s capability for complex reasoning, it comes at the cost of response efficiency.
+Finally, the s1.1 models generate considerably longer responses than the DeepSeek-R1-Distilled models.
+Despite both models being trained solely with SFT to acquire long-COT reasoning ability, the DeepSeek-R1-Distilled models use 800K training samples, while the s1.1 models are trained on only 1K.
+This discrepancy suggests that the smaller training set may lead to superficial imitation of long reasoning patterns, resulting in verbose thinking on simple questions.
+Several LRMs exhibit under-accuracy and limited robustness on simple questions.
+First, our observations find that, despite employing deep reasoning, most LRMs tend to exhibit lower accuracy on simple questions compared to traditional LLMs.
+For example, DS-R1-1.5B and EXAONE-2.4B achieve just above 50% acc@k.
+Second, many LRMs struggle with robust correctness in top-p sampling, where acc@k is significantly lower than pass@1.
+This issue is particularly pronounced in smaller LRMs.
+For instance, DS-R1-1.5B achieved 81.47% pass@1 but only 54.50% acc@k.
+5 Efficiency Analysis
+5.1 Analysis across Question Types
+To better understand the efficiency differences of LRMs across question types, we analyze the average response tokens across 4 main categories, 28 subcategories, and two languages.
+The results are displayed in Table 4 and Appendix C.2.
+LRMs exhibit a substantial increase in response length across all four major categories, 28 subcategories, and two languages.
+As shown in Table 4, for each of the four major categories, the average response length of LRMs exceeds that of LLMs by more than a factor of ten.
+Response lengths also increase significantly across all subcategories (see Appendix C.2).
+This suggests that while LRMs are primarily trained on reasoning data to produce long CoT style responses, this stylistic pattern generalizes well across a wide range of question types.
+Moreover, 23 out of the 28 LRMs produce longer responses to questions in English than Chinese.
+LRMs exhibit the most significant increase in ART for instruction following questions and tend to over-explore when the solution space is vast.
+As shown in Table 4, although small LLMs provide the most concise responses to instruction following questions, LRMs generate dramatically longer outputs—96.0 times longer in English and 57.3 times longer in Chinese than small LLMs.
+To investigate the cause, we further analyze the subcategories of instruction following questions.
+As shown in Appendix C.2, average tokens is notably longer in the subcategories of length constraints, character constraints, and sentence constraints.
+These three question types share a similar characteristic: their correctness is easy to verify, but the solution space is vast.
+We find that, although the model quickly identifies a correct answer, it becomes trapped in the search space, continually exploring alternatives and failing to stop in time.
+A case can be seen in Table 21.
+This phenomenon is more pronounced in families with lower efficiency, such as s1.1 and EXAONE.
+Figure 3: (a) Comparison of first round and additional token costs for each LRM. (b) Distribution of solution rounds for each LRM.
+5.2 Thinking Solution Analysis
+To better understand the causes of inefficiency in LRMs on S1-Bench, we analyze the solution rounds of their thinking processes [^7] We only analyze well-formatted thinking processes with correct final answers, as incorrect answers make it unclear whether LRMs are over-reasoning or under-reasoning, and malformed thinking processes cannot be precisely extracted..
+We first use DeepSeek-v3 to segment each thinking process into several solutions, each defined as a point at which LRMs explicitly arrives at a conclusion that matches the correct answer.
+We then compute the average token counts in the first solution.
+The detailed experimental setup is provided in Appendix C.3.
+Our analysis reveals the following:
+The token consumed in the first solution of LRMs significantly exceeds that of validator LLMs, as shown in Figure 3 (a).
+This suggests that LRMs may involve unnecessary reasoning steps in each solution, which could be one of the reasons for their inefficiency.
+The primary reason for efficiency gaps between LRMs lies in the number of redundant solution rounds they generate, rather than the token cost in the initial round.
+As shown in Figure 3 (a), although total thinking token counts vary widely across LRMs, their token counts in the initial round are similar and only account for a small fraction of the total.
+Figure 3 (b) further shows the distribution of solution rounds on S1-Bench, revealing that LRMs with longer thinking processes tend to generate more solution round, and this redundancy greatly increases computational cost.
+Furthermore, further experiments reveal that the redundancy in the reasoning process gradually increases over time.
+Appendix C.4 presents the experimental details.
+6 Error Analysis
+Figure 4: Distribution of the thinking process across four categories.
+FA and TP refer to Final Answer and Thinking Process, respectively.
+Green bars indicate cases where the final answer is correct, while red bars indicate cases where it is incorrect.
+This section analyzes the errors made in the thinking process.
+Specifically, we utilize DeepSeek-v3 to categorize the responses of LRMs into four cases and compute the corresponding proportions: (1) Final answer correct; thinking process entirely accurate.
+(2) Final answer correct; thinking process contains intermediate errors.
+(3) Final answer incorrect; correct answer mentioned in thinking process.
+(4) Final answer incorrect; correct answer never mentioned in thinking process.
+The classification details are in Appendix C.5; results are shown in Figure 4.
+Key findings include:
+Lower-accuracy LRMs tend to produce less reliable reasoning chains; even when they arrive at the correct final answer, their intermediate steps often contain errors (light green).
+LRMs with high accuracy (e.g., DS-R1) show almost no flawed reasoning steps, whereas those with lower accuracy (e.g., DS-R1-1.5B) often generate incorrect intermediate conclusions, further indicating that they lack robust reasoning ability.
+Although LRMs sometimes mention the correct answer during reasoning, they may deviate and ultimately produce incorrect final answers (light red).
+In one case, the LRM initially arrived at the correct answer but undermined it through excessive verification, a case can be seen in Table 24.
+In another case, the LRM directly denies the correct answer, a case can be seen in Table 23.
+7 Gut Moment
+Figure 5: Top: Count of “gut moments” across models. Bottom: Probability of “gut moments” by question type.
+[…]
+First, all LRMs show the “gut moment” phenomenon to varying degrees, which is more evident in the Qwen, DeepSeek, and Light-R1 families and Hunyuan-T1.
+Second, LRMs show stylistic differences in expressing “gut moment.“ For example, the Qwen family often views questions as simple, whereas the DeepSeek-distilled models show more diverse difficulty comments.
+Third, some LRMs show significantly stronger “gut moment” in Chinese than in English, such as the Qwen and DeepSeek families, likely due to a higher proportion of Chinese in their training data.
+Finally, the “gut moment” is most evident in reasoning questions and rarely appears in analytical questions, except in DeepSeek-distilled models.
+Figure 6: Average response tokens in the easy category vs. all samples. Dots show difference: easy minus all.
+To investigate whether the early sense of a question as “easy” leads to a corresponding reduction in response length, we compare the average response tokens for questions in the easy category versus all samples.
+The results are shown in Figure 6.
+Except for L-R1-32B, other LRMs do not exhibit a noticeable decrease in response length when questions are viewed as “easy”; in fact, 21 out of 28 LRMs showed an increase in response length under this condition.
+This suggests a discrepancy between the LRM’s initial sense of difficulty and its generative behavior, the causes and improvements of which warrant further investigation.
+8 Conclusion
+This paper introduces S1-Bench, the first benchmark designed to evaluate system 1 thinking capabilities in LRMs.
+We conduct extensive evaluations across 28 LRMs, revealing their inefficiency, inadequate accuracy, and limited robustness when handling simple questions.
+Additionally, we observe “gut moment” and find a gap between their difficulty perception and generation length.
+Overall, this work paves the way toward dual-system compatibility in the development of LRMs.
+Limitations
+Although S1-Bench pioneers the evaluation of system 1 thinking in LRMs, it still has several limitations.
+First, due to our emphasis on ensuring the uniqueness of each sample during dataset construction—for instance, including only one question for basic arithmetic operations such as addition, subtraction, and multiplication—the overall scale of the benchmark remains limited.
+As a next step, we plan to expand the scale of S1-Bench.
+Second, while recent months have seen a surge in newly released open-source LRMs, we have only evaluated 28 representative models and have not covered the full spectrum of available models.
+Lastly, we do not propose methods to improve the efficiency of LRMs on system 1 tasks in this work; this will be the focus of our future research.
+Appendix A More Information of S1-Bench Construction
+A.1 Benchmark Statistics
+We survey studies on improving the efficiency of LRMs, as there is potential overlap between these studies and the technical approaches aimed at enhancing system 1 thinking in LRMs.
+Table 7 presents the results of our survey.
+We compile the benchmarks used in these studies for evaluation, that are typically used to verify whether models achieve efficiency improvements.
+Benchmarks that appear more than four times include: MATH500 Hendrycks et al.
+(2021b), GSM8K Cobbe et al.
+(2021), AIME24/25 MAA Committees , GPQA Rein et al.
+(2024), AMC23 AI-MO (2024), MMLU Hendrycks et al.
+(2021a), Olympiad-Bench He et al.
+(2024), SVAMP Patel et al.
+(2021), LiveCodeBench Jain et al.
+(2025), and CommonSenseQA Talmor et al.
+(2019).
+The accuracy shown in Table 1 is the average result of the four models, Qwen2.5-7B, Llama3.1-8B, Mistral-8B, and Gemma2-9B, at temperature 0, using GPT-4o as the evaluator.
+A.2 Subcategories in S1-Bench
+Figure 7 shows the pie chart distribution of 28 subcategories in S1-Bench. For more details on the subcategories, please refer to Table 8,9.
+A.3 Prompt for S1-Bench construction
+This section presents the prompts used in the construction of S1-Bench, including the Initial Generation prompt, the Discriminating Generation Quality prompt, and the Reduce Difficulty prompt.
+See Table 10 for details.
+Figure 7: S1-Bench Category Display. The inner circle represents four major categories, and the outer circle includes 28 subcategories.
+Figure 8: LRMs exhibit under-accuracy and overthinking on simple problems.
+Shapes represent organizations, colors represent base model families, with darker colors indicating larger models, and connecting lines represent the relationships between model families and training.
+Appendix B Baseline Models and Evaluation Details
+B.1 Baseline Model Details
+Table 11 presents the abbreviations, IDs, and URLs of LLMs used in this paper.
+Table 12 displays the abbreviations, IDs, URLs, organizations, training algorithms, and training data volumes of open-source LRMs evaluated in this study.
+B.2 GPT-4o and Human Evaluation
+We use GPT-4o as the evaluator to assess the correctness of the responses.
+If a final answer can be isolated, only the final answer is evaluated; otherwise, the entire response is assessed.
+The evaluation prompt is provided in Table 13.
+To evaluate the consistency between the GPT-4 judge’s assessments and human judgments, we conduct a comprehensive human evaluation study involving three of the authors.
+Specifically, we randomly sample 20 question-answer pairs from each model’s greedy decoding results, resulting in a dataset of 640 pairs derived from 32 models (including 4 verifier LLMs and 28 LRMs).
+The questions, reference answers, and model responses are then presented to three annotators, who independently judge the correctness of each model response.
+The final human evaluation results are determined through majority voting.
+Ultimately, the Cohen’s Kappa between the human evaluators and the GPT-4 judge is calculated to be 0.83, indicating an exceptionally high level of agreement.
+B.3 Accuracy Metrics Details
 Pass@1: Followed DeepSeek-R1 Guo et al. (2025), we calculate pass@1 to assess the percentage of correct responses among the k=5 generations. Specifically, it is defined as:
@@
 […]
-Notably, S-Corr represents the upper bound for pass@1 and acc@5 under strict formatting requirements, with L-Corr following the same principle.
-4.3 Main Results
-Table 2 and Figure 1 presents the main results of LRMs on S1-Bench, revealing two key phenomena.
-LRMs exhibit significantly lower efficiency than LLMs on S1-Bench, and no clear correlation is observed between ART and model size.
-Notably, state-of-the-art LRMs, such as DeepSeek-R1 and QwQ-32B, do not demonstrate a distinct advantage in efficiency.
-In contrast, Sky-T1-32B, which undergoes specific optimizations to mitigate overthinking using SimPO, achieves the highest efficiency.
-The L-R1-DS 7B/14B/32B models generate longer responses than the DS-R1-7B/14B/32B models, and the former are further post-trained on the latter.
-This suggests that further post-training may improve reasoning ability for complex tasks, but at the cost of response efficiency for simple tasks.
-Finally, the s1.1 models generate considerably longer responses than the DeepSeek-R1-Distilled models.
-Despite both models being trained solely with SFT to acquire long-COT reasoning ability, the DeepSeek-R1-Distilled models use 800K training samples, while the s1.1 models are trained on only 1K.
-This discrepancy suggests that the smaller training set may lead to superficial imitation of long reasoning patterns, resulting in verbose thinking on simple questions.
-Several LRMs exhibit under-accuracy on simple questions.
-Our observations reveal that, despite employing deep reasoning, most LRMs tend to exhibit lower accuracy on simple questions compared to traditional LLMs.
-For example, DS-R1-1.5B and EXAONE-2.4B achieve just above 50% acc@k.
-Moreover, accuracy decreases with smaller model size.
-Finally, many LRMs struggle with robust correctness in top-p sampling, where acc@k is significantly lower than pass@1.
-This issue is particularly pronounced in smaller LRMs.
-For instance, DS-R1-1.5B achieved 81.47% pass@1 but only 54.50% acc@k.
-5 Efficiency Analysis
-Figure 4: (a) Comparison of Initial and Additional Thinking Costs for Each LRM. (b) Distribution of Solution Rounds for Each LRM.
-5.1 ART Analysis of Question Types
-To better understand the efficiency differences of LRMs across question types, we analyze $\mathtt{ART}$ across 4 main categories, 28 subcategories, and two languages, yielding the following insights.
-LRMs exhibit the most significant increase in ART for instruction-following questions and tend to over-explore when the solution space is vast.
-Although small LLMs generate the most concise responses for instruction-following questions across all main categories in S1-Bench, LRMs produce responses that are 109.7× longer in Chinese and 62.7× longer in English, making them the longest response type across all categories.
-We further analyze the subcategories of instruction-following questions.
-As shown in Figure 9, $\mathtt{ART}$ is notably longer in the subcategories of length constraints, character constraints, and sentence constraints.
-These three question types share a similar characteristic: their correctness is verifiable, but the solution space is vast.
-We find that, although the model quickly identifies a correct answer, it becomes trapped in the search space, continually exploring alternatives and failing to stop in time.
-This phenomenon is more pronounced in families with higher $\mathtt{ART}$, such as s1.1 and the EXAONE.
-LRMs also demonstrate a significant increase in ART for reasoning tasks.
-In Table 3, $\mathtt{ART}$ for reasoning questions also increases significantly, and Figure 9 further corroborates this phenomenon in $\mathtt{ART}$ in various reasoning subcategories.
-One possible explanation for this is that reasoning questions align closely with the post-training data distribution of LRMs, thereby further stimulating the long-chain reasoning behavior of LRMs.
-LRMs consistently exhibit inefficiency in both Chinese and English.
-In table 3, most LRMs show higher $\mathtt{ART}$ on English questions, while QwQ-32B and Hunyuan-T1 exhibit comparable $\mathtt{ART}$ across Chinese and English.
-Figure 5: Maximum similarity between each segment and all preceding segments for LRMs across four categories.
-5.2 Solution Analysis in Thinking Processes
-To better understand the causes of inefficiency in LRMs on the S1-Bench, we analyze the thinking process where the final answer is correct and the format is strictly correct and non-empty [^9] We selected this subset because incorrect answers make it unclear whether LRMs are over-reasoning or under-reasoning, and malformed or empty thinking processes cannot be precisely extracted..
-We begin by segmenting each thinking process into several solutions, each defined as a point at which LRMs explicitly arrive at a conclusion that directly aligns with the correct answer.
-The segmentation process is performed by DeepSeek-v3, with prompts detailed in Table 12.
-We then compute the average initial thinking cost for LRMs.
-For each sample, if the thinking process contains at least one solution, the cost is defined as the tokens counts in the first solution.
-If no clear and correct solution is provided, the cost is the total token counts in the thinking process.
-The reported value is the average across all selected samples.
-We find that:
-The step-by-step reasoning in the initial thinking process is not the primary cause of the inefficiency in LRMs.
-As shown in Figure 4 (a), although the total counts of thinking tokens varies considerably across LRMs, their initial thinking costs are similar and account for only a small fraction of the total.
-Generating unnecessary solution rounds after reaching the correct answer is one of the reasons for the inefficiency of LRMs.
-We further examine the distribution of solution rounds among various LRMs on the S1-Bench (Figure 4 (b)) and find that models with longer thinking processes tend to produce excessive solution rounds, repeatedly re-verifying simple problems that have already been solved.
-This redundant verification significantly contributes to computational inefficiency.
-Figure 6: Distribution of the thinking process across four categories.
-FA and TP refer to Final Answer and Thinking Process, respectively.
-Green bars indicate cases where the final answer is correct, while red bars indicate cases where it is incorrect.
-5.3 Redundancy in Thinking Processes
-Information redundancy increases as reasoning sequences increase.
-We conduct a similarity analysis to analyze how information redundancy in the thinking processes of LRMs changes as reasoning sequences increase.
-Specifically, we first divide the complete thinking process into k equal-length segments [^10] We set k=15, changing its value does not affect the conclusions..
-Then, we encode each segment using the all-MiniLM-L6-v2 model [^11] https://huggingface.co/sentence-transformers/allMiniLM-L6-v2.
-For each segment, we calculate the cosine similarity with all its preceding segments and use the maximum similarity as a measure of its information redundancy.
-As shown in Figure 5, information redundancy increases across all four main categories as reasoning sequences increase.
-Sky-T1-32B shows overall lower similarity, which stems from its shorter thinking process, but still demonstrates an upward trend.
-6 Analysis of Errors
-Figure 7 explores the relationship between thinking process quality and final answer accuracy.
-We focus on samples whose format is strictly correct and thinking process is non-empty.
-For samples with correct final answers, we categorize them based on whether the thinking process contains explicit incorrect conclusions in intermediate steps.
-For samples with incorrect final answers, we categorize them based on whether the correct answer is mentioned at least once during reasoning.
-We use DeepSeek-v3 for categorization, with the prompt provided in Table 13.
-Figure 6 presents the distribution of these four categories in different RLMs.
-Our analysis yields the following observations: (1) LRMs with lower accuracy often include incorrect intermediate conclusions in their reasoning, even when they ultimately reach correct final answers (light green).
-(2) Although LRMs sometimes reach the correct answer during reasoning, they may deviate and ultimately produce incorrect conclusions (light red).
-7 Simplicity Prejudgement
-Figure 7: Counting instances of prejudgment generated by LRMs for simple questions and $\mathtt{ART}$.
-Finally, we discover an intriguing phenomenon: LRMs can prejudge certain simple questions.
-Specifically, we extract the first 50 tokens from each thinking process and use DeepSeek-v3 to identify sentences showing judgment, further extracting judgment words [^12] An example of prejudgment: “This question seems simple”.
-For prompt references, see Table 14..
-Figure 7 displays the frequency of words judging questions as simple, with details provided in the Table 7.
-Our findings reveal:
-LRMs possess the ability to prejudge question simplicity, especially in Chinese.
-All LRMs exhibit prejudgment phenomena within their thinking processes, demonstrating an ability to directly assess question difficulty.
-Additionally, 14 LRMs exhibit significantly stronger prejudgment tendencies in Chinese, with this phenomenon concentrating primarily within the L-R1, DS-R1, QwQ, and Hunyuan model families.
-Even with prejudgment, the thinking length of LRMs does not shorten.
-As shown in Figure 7, the average $\mathtt{ART}$ of the thinking process exhibiting prejudgment does not decrease.
-We consider further exploration of this phenomenon as an important direction for future work.
-These results suggest LRMs possess an inherent understanding of question difficulty. This opens a novel pathway toward dual-system compatibility for LRMs, which we identify as future work.
-8 Conclusion
-This paper presents S1-Bench, designed to evaluate LRMs’ performance on simple questions suited for system 1 thinking.
-Experiments reveal significant overthinking and performance degradation by LRMs on S1-Bench, highlighting the gap between current LRMs and dual-system compatibility.
-Appendix A Subcategories in S1-Bench
-Figure 8: S1-Bench Category Display. The inner circle represents four major categories, and the outer circle includes 28 subcategories.
-Figure 8 shows the pie chart distribution of 28 subcategories in S1-Bench. For more details on the subcategories, please refer to Table 15,16.
-Appendix B Baseline Model Details
-Model Model ID URL
-Qwen2.5-7B Qwen2.5-7B-Instruct https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
-Llama3.1-8B Llama-3.1-8B-Instruct https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
-Mistral-8B Ministral-8B-Instruct-2410 https://huggingface.co/mistralai/Ministral-8B-Instruct-2410
-Gemma2-9B gemma-2-9b-it https://huggingface.co/google/gemma-2-9b-it
-Qwen2.5-14B Qwen2.5-14B-Instruct https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
-Qwen2.5-32B Qwen2.5-32B-Instruct https://huggingface.co/Qwen/Qwen2.5-32B-Instruct
-Qwen2.5-72B Qwen2.5-72B-Instruct https://huggingface.co/Qwen/Qwen2.5-72B-Instruct
-Llama3.3-70B Llama-3.3-70B-Instruct https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct
-DeepSeek-v3 DeepSeek-V3-0324 https://huggingface.co/deepseek-ai/DeepSeek-V3-0324
-Table 4: Mapping of LLM abbreviations and IDs used in this paper, with their open-source URLs.
-Model ID Abbreviation Base Model Alg. Size
-DeepSeek
-DeepSeek-R1-Distill-Qwen-1.5B DS-R1-1.5B Qwen2.5-Math-1.5B SFT 800K
-DeepSeek-R1-Distill-Qwen-7B DS-R1-7B Qwen2.5-Math-7B SFT 800K
-DeepSeek-R1-Distill-Llama-8B DS-R1-8B Llama-3.1-8B SFT 800K
-DeepSeek-R1-Distill-Qwen-14B DS-R1-14B Qwen2.5-14B SFT 800K
-DeepSeek-R1-Distill-Qwen-32B DS-R1-32B Qwen2.5-32B SFT 800K
-DeepSeek-R1-Distill-Llama-70B DS-R1-70B Llama-3.3-70B-Instruct SFT 800K
-DeepSeek-R1 DS-R1 DeepSeek-V3-0324 SFT&RL 800K&–
-Qwen
-QwQ-32B QwQ-32B Qwen2.5-32B
-qihoo360
-Light-R1-7B-DS L-R1-7B-DS DeepSeek-R1-Distill-Qwen-7B SFT 3K
-Light-R1-14B-DS L-R1-14B-DS DeepSeek-R1-Distill-Qwen-14B SFT&RL 3K&–
-Light-R1-32B-DS L-R1-32B-DS DeepSeek-R1-Distill-Qwen-32B SFT 3K
-Light-R1-32B L-R1-32B Qwen2.5-32B-Instruct SFT&DPO 73K&–
-simplescaling
-s1.1-7B s1.1-7B Qwen2.5-7B-Instruct SFT 1K
-s1.1-14B s1.1-14B Qwen2.5-14B-Instruct SFT 1K
-s1.1-32B s1.1-32B Qwen2.5-32B-Instruct SFT 1K
-LG AI Research
-EXAONE-Deep-2.4B EXAONE-2.4B EXAONE-3.5-2.4B-Instruct SFT&DPO&RL 1.6M&20K&10K
-EXAONE-Deep-7.8B EXAONE-7.8B EXAONE-3.5-7.8B-Instruct SFT&DPO&RL 1.6M&20K&10K
-EXAONE-Deep-32B EXAONE-32B EXAONE-3.5-32B-Instruct SFT&DPO&RL 1.6M&20K&10K
-NVIDIA
-Llama-3.1-Nemotron-Nano-8B-v1 Nemotron-8B Llama-3.1-8B-Instruct SFT&RL –&–
-Llama-3.3-Nemotron-Super-49B-v1 Nemotron-49B Llama-3.3-70B-Instruct SFT&RL –&–
-NovaSky
-Sky-T1-32B-Flash Sky-T1-32B Qwen2.5-32B-Instruct SFT&SimPO 17K&10K
-Table 5: The open-source LRMs details evaluated for S1-Bench.
-Table 4 presents the abbreviations, IDs, and URLs of LLMs used in this paper.
-Table 5 displays the abbreviations, IDs, URLs, organizations, algorithms, and training data volumes of open-source LRMs evaluated in this study.
-Appendix C Format Details
+B.4 Types and Analysis of Format Errors
+This section introduces a comprehensive taxonomy of format errors and highlights the importance of addressing these issues in future research.
+Unlike conventional LLMs, LRMs frequently exhibit format errors.
+These errors are defined by failing to use a unique end thinking marker (e.g., </think>) to separate the thinking process from the final answer.
+Format errors increase the difficulty of distinguishing the thinking process from the final answer and reveal the vulnerability of LRMs in following predefined formats.
+To illustrate this phenomenon, we identify 12 distinct types of response formats produced by LRMs, each assigned a unique ID, as shown in Table 5.
+These 12 types are further grouped into three major categories:
+  • Standard-Conforming Responses: These responses meet the expected format by including exactly one end thinking marker (e.g., </think>) to delimit the thinking process from the final answer.
+Among these, type ID-100 includes a thinking process, while ID-101 omits it.
+The proportion of such responses is measured using the S-Corr metric.
+  • Unreadable Responses: These refer to generation failures, including cases where LRMs produce endlessly thinking content or solely produce end thinking markers.
+The proportion of all other (i.e., readable) responses is measured using the L-Corr metric.
+[…]
+Table 14 and Table 15 present the distributions of 12 format types under top-p sampling and greedy sampling, respectively.
+we find: (1) The infinite generation phenomenon is widespread across most LRMs, particularly concentrated in LRMs with fewer than 32B parameters.
+(2) The Nemotron and EXAONE families frequently produce correctly formatted responses without any explicit thinking processes.
+This behavior can be viewed as a mechanism for mitigating over-thinking.
+However, the EXAONE family still exhibits substantial overthinking tendencies, suggesting that LRMs’ capability to respond without visible reasoning and their tendency to overthink may be orthogonal characteristics.
+(3) None of the evaluated LRMs exhibited behaviors classified as ID-205/206.
 Format ID marker marker marker thinking final
-(standard) (special) (number) process answer
-Strict 100 $\surd$ 1 $\surd$ $\surd$
+(standard) (unmatched) (number) process answer
+Standard 100 $\surd$ 1 $\surd$ $\surd$
 101 $\surd$ 1 $\times$ $\surd$
-Loose (w/o Strict) 200 $\surd$ 1 $\surd$ $\times$
+Readable but Malformed 200 $\surd$ 1 $\surd$ $\times$
 201 $\surd$ >1 $\surd$ $\surd$
@@
 207 $\times$ $\times$ 0 $\surd$
-Error 300 $\surd$ $\surd$ $\geq$1 $\times$ $\times$
+Unreadable 300 $\surd$ $\surd$ $\geq$1 $\times$ $\times$
 301 $\times$ $\times$ 0 $\infty$
-Table 6: Format types.
-Words Count
-{CJK}UTF8gkai简单 466
-straightforward 231
-simple 118
-{CJK}UTF8gkai不难 72
-{CJK}UTF8gkai基础的 44
-basic 24
-{CJK}UTF8gkai直接 15
-{CJK}UTF8gkai常见 10
-{CJK}UTF8gkai基础 10
-{CJK}UTF8gkai不复杂 7
-simplest 7
-pretty straightforward 5
-easy 5
-{CJK}UTF8gkai基本的 5
-{CJK}UTF8gkai基本 3
-Table 7: Top 15 phrases with prejudgment phenomenon, with counts summed across 22 LRMs.
-Model Loose Format Error Format
-Strict Format
-100 101 200 201 202 203 204 205 206 207 300 301
-QwQ-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-hunyuan-t1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-Sky-T1-32B 95.26 0.00 0.62 0.09 0.19 0.00 0.28 0.00 0.00 3.03 0.00 0.52
-DS-R1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-DS-R1-70B 99.91 0.00 0.09 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-DS-R1-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-DS-R1-14B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-DS-R1-8B 99.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24 0.00 0.24
-DS-R1-7B 99.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.76
-DS-R1-1.5B 97.58 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.42
-Nemotron-49B 66.07 33.93 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-Nemotron-8B 58.06 26.26 0.00 0.00 0.00 0.09 0.00 0.00 0.00 15.02 0.00 0.57
-L-R1-32B 95.07 0.00 0.00 0.00 0.00 0.00 0.81 0.00 0.00 3.03 0.00 1.09
-L-R1-32B-DS 99.81 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.19
-L-R1-14B-DS 99.19 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.81
-L-R1-7B-DS 99.67 0.00 0.05 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
-s1.1-32B 99.53 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.43
-s1.1-14B 97.39 0.00 0.00 0.14 0.00 0.00 0.24 0.00 0.00 0.00 0.00 2.23
-s1.1-7B 88.96 0.00 0.00 7.96 0.09 0.00 0.00 0.00 0.00 0.09 0.00 2.89
-EXAONE-32B 67.39 32.42 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.19
-EXAONE-7.8B 65.83 32.23 0.00 0.00 0.05 0.47 0.00 0.00 0.00 0.14 0.00 1.28
-EXAONE-2.4B 81.42 15.83 0.00 0.09 0.00 0.05 0.00 0.00 0.00 0.05 0.00 2.56
-Table 8: Format types rates in top-p sampling setting.
-Model Loose Format Error Format
-Strict Format
-100 101 200 201 202 203 204 205 206 207 300 301
-QwQ-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-hunyuan-t1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-Sky-T1-32B 99.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.47 0.00 0.24
-DS-R1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-DS-R1-70B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-DS-R1-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-DS-R1-14B 99.76 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
-DS-R1-8B 99.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24 0.00 0.24
-DS-R1-7B 97.87 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.13
-DS-R1-1.5B 91.94 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 8.06
-Nemotron-49B 60.90 39.10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-Nemotron-8B 55.21 26.78 0.00 0.00 0.00 0.00 0.00 0.00 0.00 16.35 0.00 1.66
-L-R1-32B 85.55 0.24 0.00 0.24 0.71 0.24 0.95 0.00 0.00 6.64 2.61 2.84
-L-R1-32B-DS 99.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.71
-L-R1-14B-DS 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
-L-R1-7B-DS 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
-s1.1-32B 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
-s1.1-14B 95.97 0.00 0.00 0.24 0.00 0.00 0.24 0.00 0.00 0.00 0.00 3.55
-s1.1-7B 87.91 0.00 0.00 6.64 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.45
-EXAONE-32B 65.88 33.89 0.00 0.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
-EXAONE-7.8B 63.51 33.65 0.00 0.00 0.00 0.24 0.00 0.00 0.00 0.24 0.00 2.37
-EXAONE-2.4B 78.91 15.88 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.21
-Table 9: Format types rates in greedy decoding setting.
-C.1 Division of Different Format
-This section presents a comprehensive taxonomy of format errors and emphasizes the importance of addressing these issues in future research endeavors.
-Unlike conventional LLMs, LRMs frequently exhibit format errors, manifested when responses fail to adhere to the requirement of generating a single End Thinking Marker ($\mathtt{ETM}$).
-This inconsistency creates significant challenges in distinguishing between the thinking process and the final answer.
-Format errors represent a distinct category from over-thinking and under-accuracy phenomena, yet similarly indicate vulnerability to elementary reasoning traps.
-As illustrated in Table 6, we have identified and classified 12 distinct format types, each assigned a unique ID.
-These format types can be broadly categorized into Strict Format and Loose Format, with Strict Format representing a specialized subset of Loose Format.
-Loose Format.
-ID-100 and ID-101 represent two standard decoding modes, categorized as Strict Format.
-Both contain only one $\mathtt{ETM}$.
-Among these, ID-100 includes a thinking process, while ID-101 only generates the $\mathtt{ETM}$ and final answer.
+Table 5: Twelve types of response format.
+Appendix C More Experimental Setups & Results
+C.1 Greedy Sampling Results
+Table 16 presents the performance of LRMs on S1-Bench under greedy sampling.
+While overall accuracy improves compared to top-p sampling, issues of inefficiency and accuracy degradation on simple questions remain.
+C.2 Efficiency Analysis across Subcategories.
+Figure 9 illustrates the average response tokens across the 28 subcategories.
+In the heatmap, both models (rows) and subcategories (columns) are ordered in descending order according to their average number of response tokens.
+C.3 Solution Analysis Details
+For solution analysis, We only use well-formatted thinking processes with correct final answers, as incorrect answers make it unclear whether LRMs are over-reasoning or under-reasoning, and malformed thinking processes cannot be precisely extracted.
+The segmentation process is performed by DeepSeek-v3, with prompts detailed in Table 17.
+We compute the average token count in the first solution round; if no solution is found, we use the token count of the entire thinking process.
+C.4 Thinking Redundancy Analysis
+We conduct a similarity analysis to analyze how information redundancy in the thinking processes changes as reasoning sequences increase.
+Specifically, we first divide the complete thinking process into k equal-length segments [^10] We set k=15, changing its value does not affect the conclusions..
+Then, we encode each segment using the all-MiniLM-L6-v2 model [^11] https://huggingface.co/sentence-transformers/allMiniLM-L6-v2.
+For each segment, we calculate the cosine similarity with all its preceding segments and use the maximum similarity as a measure of its information redundancy.
+As shown in Figure 10, information redundancy increases across all four main categories as reasoning sequences increase.
+Sky-T1-32B shows overall lower similarity, which stems from its shorter thinking process, but still demonstrates an upward trend.
+C.5 Error Analysis Details
+In error analysis, we only use well-formatted samples, as malformed thinking processes cannot be precisely extracted.
+For samples with correct final answers, we categorize them based on whether the thinking process contains explicit incorrect conclusions in intermediate steps.
+For samples with incorrect final answers, we categorize them based on whether the correct answer is mentioned at least once during reasoning.
+We use DeepSeek-v3 for categorization, with prompts provided in Table 18.
+C.6 Gut Moment Analysis Details
 […]
-Error Format.
-LRMs produce two types of undecodable errors: ID-300 indicates that the LRM outputs only a $\mathtt{ETM}$ without any additional content, while ID-301 represents cases where the LRM’s thinking process reaches the maximum length.
-C.2 Evaluation of Different Format
-In the classification shown in Table 6, ID-100 / 101 / 201 / 203 / 204 / 206 / 207 have decomposable final answers.
-If multiple $\mathtt{ETM}$s appear, only the last one is selected for splitting.
-In this case, the evaluator uses the final answer as the evaluation target.
-ID-200 / 202 / 205 lack a final answer, in which case the evaluation target becomes the thinking process, as this part may also contain the correct solution to the question.
-C.3 Statistics of Different Format
-Table 8 presents the distribution of 12 format types under top-p sampling settings.
-Our analysis reveals three key insights: (1) The infinite thinking phenomenon is widespread across most models, particularly concentrated in LRMs with fewer than 32B parameters.
-This indicates that output formatting should be prioritized during training to reduce format errors, especially when working with less powerful foundation models.
-(2) The Nemotron and EXAONE families frequently produce correctly formatted responses without visible reasoning processes.
-This behavior can be viewed as a mechanism for mitigating over-thinking.
-However, the EXAONE family still exhibits substantial overthinking tendencies, suggesting that LRMs’ capability to respond without visible reasoning and their tendency to overthink may be orthogonal characteristics.
-(3) None of the evaluated LRMs exhibited behaviors classified as ID-205/206, possibly because special $\mathtt{ETM}$s are perceived by LRMs as the only alternative to standard $\mathtt{ETM}$s.
-Furthermore, ID-300 was not observed, indicating that high-temperature sampling generally yields semantically meaningful responses.
-Table 9 shows the format statistics of greedy sampling.
-The phenomenon is generally similar to top-p sampling.
-L-R1-32B produced ID-300, possibly due to poor compatibility between the training method and low temperature.
-Appendix D Experiment Details
-Table 10 shows LRMs’ performance on S1-Bench under greedy sampling settings.
-While overall accuracy improves compared to top-p sampling settings, $\mathtt{ART}$ slightly decreases, yet overthinking and accuracy degradation issues persist.
-Figure 9 illustrates the $\mathtt{ART}$ heatmap across 28 subcategories and 26 models.
-We compile the statistics of the prejudgment words extracted, as shown in Table 7.
-Model Size acc (Loose) acc (Strict) $\mathtt{SUC}$-$\mathtt{L}\uparrow$ $\mathtt{SUC}$-$\mathtt{S}\uparrow$ $\mathtt{ART}$downarrow
-QwQ-32B 32B 100.00 100.00 100.00 100.00 750.41
-Hunyuan-T1 100.00 100.00 100.00 100.00 541.09
-Sky-T1-32B 32B 99.53 99.05 99.76 99.29 157.12
-DS-R1 671B 100.00 100.00 100.00 100.00 621.89
-DS-R1-70B 70B 99.76 99.76 100.00 100.00 469.78
-DS-R1-32B 32B 100.00 100.00 100.00 100.00 428.46
-DS-R1-14B 14B 99.29 99.29 99.76 99.76 463.52
-DS-R1-8B 8B 97.63 97.39 99.76 99.53 452.11
-DS-R1-7B 7B 94.31 94.31 97.87 97.87 436.87
-DS-R1-1.5B 1.5B 76.54 76.54 91.94 91.94 473.67
-Nemotron-49B 49B 99.53 99.53 100.00 100.00 337.94
-Nemotron-8B 8B 84.60 77.73 98.34 81.99 446.62
-L-R1-32B 32B 92.18 85.78 94.55 85.78 996.36
-L-R1-32B-DS 32B 99.29 99.29 99.29 99.29 528.45
-L-R1-14B-DS 14B 98.82 98.82 98.82 98.82 664.28
-L-R1-7B-DS 7B 92.65 92.65 98.82 98.82 514.60
-s1.1-32B 32B 98.82 98.82 98.82 98.82 983.38
-s1.1-14B 14B 95.97 95.50 96.45 95.97 786.30
-s1.1-7B 7B 94.31 87.68 94.55 87.91 630.52
-EXAONE-32B 32B 97.63 97.39 100.00 99.76 746.89
-EXAONE-7.8B 7.8B 86.73 86.49 97.63 97.16 947.92
-EXAONE-2.4B 2.4B 72.99 72.99 94.79 94.79 1394.72
-Table 10: Main results in the greedy decoding setting on the S1-Bench, sorted by model family. Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst.
-Figure 9: $\mathtt{ART}$ on the 28 subcategories, which is the average result of five generations under top-p sampling with Loose Format setting.
-Appendix E Prompts discussed in this paper
-This section presents the prompts used in this paper. For S1-Bench construction-related prompts, please refer to Table 11; for experimental prompts, please refer to Table 12,13 and 14.
-Prompt for construction workflow for S1-Bench
-Initial Generation
-Based on the category name and explanation, generate 50 pairs of questions and answers in both Chinese and English.
-Ensure that the generated questions vary in perspective and content, maintaining diversity.
-All questions must have a **uniquely determined, unambiguous answer with no potential conditions causing multiple possible answers**.
-Ensure that the questions are expressed clearly.
-Additionally, the difficulty level of the questions should be **as simple as possible—so easy that they can be answered correctly without any need for thought**.
-# Category Name:
-{category}
-# Explanation:
-{explanation}
-# Cases:
-## English question:
-{question_en}
-## English Answer:
-{answer_en}
-## Chinese question:
-{question_zh}
-## Chinese Answer:
-{answer_zh}
-Please generate 50 pairs of Chinese and English questions and answers in the following format: [question]English-question[answer]English-answer[question]Chinese-question[answer]Chinese-answer…
-Start generating:
-Discriminating Generation Quality
-Given a question, its answer, and its category, please analyze from the following perspectives as comprehensively as possible:
-1. Whether the question belongs to the specified category.
-2. Whether the question is clear, unambiguous, has an absolutely unique answer, and has no potential alternative interpretations.
-3. Whether the answer is absolutely correct; if not, what the correct answer should be.
-4. Whether the question is similar to other given questions, and if similar, whether more diverse questions can be generated.
-# Question Category and Explanation:
-{category_with_exp}
-# Question and Answer:
-{question_with_answer}
-# Other Questions:
-{questions_list}
-Begin your analysis, aiming to be as detailed and comprehensive as possible:
-Reduce Difficulty Generation
-Given a question and answer that are too complex for the model to answer correctly, you need to further reduce their difficulty while trying to:
-- Maintain the original perspective and semantic information
-- Preserve the question’s alignment with its category
-# Question Category and Explanation:
-{category_with_exp}
-# Question and Answer:
-{question_with_answer}
-The new question and answer:
-Evaluation on S1-Bench
-**Question:**
-{question}
-**Model Answer:**
-{model_answer}
-**Ground Truth:**
-{ground_truth}
-Your task is to evaluate whether the model’s answer is correct.
-An answer is considered correct as long as it contains the ground truth (regardless of how complex or detailed the description is).
-If there are parenthetical notes after the ground truth, then there may be multiple correct answers.
-In this case, the given answer is just one example, and any answer that meets the requirements specified in the notes can be considered correct.
-Additionally, some reasonably uncertain supplementary information is also considered appropriate, including more details, possibilities, and expanded discussion.
-You should focus more on whether the reply contains the correct answer.
-You need to output a standard JSON, providing your explanation of the evaluation in the "explain" field, and giving the evaluation result in the "result" field, where 1 means the answer is correct and 0 means it is incorrect.
-Your action should follow the given format: "explain": "", "result": 0/1
-Table 11: Prompt for construction workflow for S1-Bench.
-The “explanation” (and “category_with_exp”) includes the definition of each major category, simplicity requirements in Section 3.1.1, and the definition of subcategories.
-Prompts for Solution Segmentation
-# Task Description:
-Your task is to segment the given Chain of Thought according to the following rules:
-1. Segmentation positions:
-(1) Please identify and extract all sub-reasoning processes from the Chain of Thought that meet the following condition: They explicitly arrive at a conclusion (including cases phrased as questions, e.g., "right?") that is directly consistent with the Ground Truth.
-Reasoning processes that only indirectly support the Ground Truth or result in partially aligned conclusions should be excluded.
-(2) After clearly reaching the conclusion consistent with the Ground Truth, insert the segmentation marker <split> to distinctly separate each qualifying sub-reasoning process.
-2. Output Restriction:
-(1) You should only directly output the segmentation result without adding any additional supplements.
-(2)Except for inserting the <split> separator, you must not make any other modifications to the original Chain of Thought, not even minor character-level changes such as punctuation, spacing, or capitalization.
-In other words, after removing all <split> separators, the output text must match the original Chain of Thought exactly.
-# Some cases:
-## Case 1:
-### Question:
-What is the value of x if x plus 3 equals 8?
-### Ground Truth:
-5
-### Chain of Thought:
-Okay, so I need to figure out what x is if x plus 3 equals 8. The problem says "x plus 3 equals 8," so in mathematical terms, that would be: x + 3 = 8
-Alright, now I need to solve for x.
-From what I remember in algebra, the goal is to isolate x on one side of the equation.
-That means I need to get rid of that "+ 3" that’s being added to x.
-To do that, I should perform the opposite operation.
-Since it’s adding 3, the opposite would be subtracting 3.
-So, if I subtract 3 from both sides, that should keep things equal. Let me try that: x + 3 - 3 = 8 - 3
-Simplifying both sides, on the left side, the +3 and -3 cancel each other out, leaving just x. And on the right side, 8 minus 3 is 5. So that simplifies to: x = 5
-Wait, let me check if that makes sense. If I substitute x back into the original equation, does it hold true? Let’s see:
-Original equation: x + 3 = 8
-Substitute x = 5: 5 + 3 = 8
-5 plus 3 is indeed 8, so that works out.
-Another way might be to consider what number, when you add 3 to it, gives 8.
-So starting from 8, if I take away 3, I should get back to the original number.
-That’s the same as subtracting 3 from 8, which again gives 5.
-So that’s consistent with the first method.
-Wait a second, what if the original problem was written differently? The problem says "x plus 3 equals 8." In symbols, that’s unambiguous: x + 3 = 8. So no, I think that’s right.
-*Final Answer**
-The value of x is 555.
-### Output:
-Okay, so I need to figure out what x is if x plus 3 equals 8. The problem says "x plus 3 equals 8," so in mathematical terms, that would be:x + 3 = 8
-Alright, now I need to solve for x.
-From what I remember in algebra, the goal is to isolate x on one side of the equation.
-That means I need to get rid of that "+ 3" that’s being added to x.
-To do that, I should perform the opposite operation.
-Since it’s adding 3, the opposite would be subtracting 3.
-So, if I subtract 3 from both sides, that should keep things equal. Let me try that: x + 3 - 3 = 8 - 3
-Simplifying both sides, on the left side, the +3 and -3 cancel each other out, leaving just x. And on the right side, 8 minus 3 is 5. So that simplifies to: x = 5<split>
-Wait, let me check if that makes sense. If I substitute x back into the original equation, does it hold true? Let’s see:
-Original equation: x + 3 = 8
-Substitute x = 5: 5 + 3 = 8
-5 plus 3 is indeed 8, so that works out.<split>
-Another way might be to consider what number, when you add 3 to it, gives 8.
-So starting from 8, if I take away 3, I should get back to the original number.
-That’s the same as subtracting 3 from 8, which again gives 5.<split> So that’s consistent with the first method.
-Wait a second, what if the original problem was written differently? The problem says "x plus 3 equals 8." In symbols, that’s unambiguous: x + 3 = 8. So no, I think that’s right.
-*Final Answer**
-The value of x is 555.<split>
-## Case 2:
-## Case 3:
-# Eval Target:
-## Question:
-{question}
-## Ground Truth:
-{ground_truth}
-## Chain of Thought:
-{thinking_process}
-## Output:
-Table 12: Prompts for Solution Segmentation.
-Prompts for Error Analysis
-Types of Correct
-# Task Description:
-You will receive a Question, its corresponding Ground Truth, and a Chain of Thought(COT) generated by a LLM for that Question.
-Your task is to carefully analyze the CoT and assign it to one of the two predefined categories listed below.
-# Categories:
-1: The CoT ***includes explicit incorrect conclusions*** in intermediate reasoning steps.
-2: The CoT ***doesn’t include any explicit incorrect conclusion*** in intermediate reasoning steps.
-Output your evaluation in the following format:
-# TheReason: [note: Conduct a step-by-step analysis, stating if and where explicit incorrect conclusions occur in the COT.]
-# ErrorType: [note: Summarize each incorrect conclusion into a specific error type using a phrase of less than 5 words, such as factual inaccuracies, logical fallacies, comprehension mistakes, calculation errors, formatting issues, and so forth, to better conduct further evaluation and analysis.
-Directly output a Python list, where each element represents the error type of a specific incorrect conclusion in the CoT.
-If there are no incorrect conclusions, return an empty list.]
-# TheCategory: [note: Provide your classification based on your analysis using only the number "1" or "2". Do not add any additional text.]
-# Question:
-{question}
-# Ground Truth:
-{gound_truth}
-# COT:
-{thinking_process}
-# TheReason:
-Types of Error
-# Task Description:
-You will receive a Question, its corresponding Ground Truth, and a Chain of Thought(COT) generated by a LLM for that Question.
-Your task is to carefully analyze the CoT and assign it to one of the two predefined categories listed below.
-# Categories:
-1: Regardless of whether the CoT ultimately arrives at the correct final answer or not, ***the correct answer is explicitly mentioned at least once*** within the reasoning steps (even if it is not ultimately adopted).
-2: ***The correct answer is never explicitly mentioned or referenced*** at any point within the reasoning steps.
-Output your evaluation in the following format:
-# TheReason: [note: Conduct a step-by-step analysis, explicitly stating whether and where a correct answer is mentioned within the reasoning steps.]
-# TheCategory: [note: Provide your classification based on your analysis using only the number "1" or "2". Do not add any additional text.]
-# Question:
-{question}
-# Ground Truth:
-{answer}
-# COT:
-{thinking_part}
-# TheReason:
-Table 13: Prompts for Error Analysis.
-Prompts for Extracting Prejudgment Sentences and Words
-Extracting Sentences
-# Given the beginning of a model’s thought process when answering a question, determine if there is any evaluation by the model regarding the difficulty level or category of the question.
-# For example:
-Example Evaluation 1: This seems to be an easy question.
-Example Evaluation 2: {CJK}UTF8gkai这看起来似乎是个基础的问题
-Example Evaluation 3: This question is a bit complex.
-Example Evaluation 4: {CJK}UTF8gkai这个问题看起来复杂
-# Beginning of thought process:
-{thinking_process}
-# Please output only the evaluative short phrase you find. If there is no evaluation, output "None".
-# Output:
-Extracting Words
-# You are given a comment sentence, and your goal is to find the only comment-related word.
-# For example:
-## case 1:
-Sentence: {CJK}UTF8gkai这是一个很简单的算术题
-Word: {CJK}UTF8gkai简单
-## case 2:
-Sentence: that seems pretty straightforward
-Word: straightforward
-# The comment sentence:
-{review_sentence}
-Output the only comment-related word:
-Table 14: Prompts for Extracting Prejudgment Sentences and Words.
+Type Sentence Count
+easy-zh {CJK}UTF8gkai 这个问题看起来挺简单的 308
+easy-en that seems straightforward 36
+difficult-zh {CJK}UTF8gkai 这个问题看起来有点复杂 308
+difficult-en percentages can sometimes be tricky 7
+neutral-zh {CJK}UTF8gkai 这个问题看起来好像不难 24
+neutral-en Hmm, interesting 3
+Table 6: The most common sentence in each type of “gut moment."
+Appendix D Error Cases
+This section presents several error cases observed in LRMs. See Tables 2021222324, and 25.
+Paper Abbreviation MATH GSM8K AIME GPQA AMC MMLU Olympiad- SVAMP LiveCode- Common-
+Bench Bench SenseQA
+Codi Shen et al. (2025b)
+CISC Taubenfeld et al. (2025)
+CoT-Valve Ma et al. (2025b)
+Dast Shen et al. (2025a)
+ATM Chen et al. (2024a)
+DEER Yang et al. (2025b)
+DPTS Ding et al. (2025)
+Dynasor Fu et al. (2024)
+ESC Li et al. (2024)
+Hawkeye She et al. (2025)
+token complexity Lee et al. (2025)
+INFTYTHINK Yan et al. (2025)
+KIMI K1.5 Team et al. (2025a)
+L1 Aggarwal and Welleck (2025)
+LightThinker Zhang et al. (2025a)
+LS-Mixture SFT Yu et al. (2025a)
+DSC Wang et al. (2024)
+O1-Pruner Luo et al. (2025)
+MRT Qu et al. (2025b)
+Self-Doubt Fu et al. (2025)
+RASC Wan et al. (2024)
+NoThinking Ma et al. (2025a)
+Retro-Search Lu et al. (2025)
+RSD Liao et al. (2025)
+ST-BoN Wang et al. (2025b)
+Elastic Reasoning Xu et al. (2025b)
+FS-BoN Munkhbat et al. (2025)
+SoT Aytes et al. (2025)
+SpecReason Pan et al. (2025)
+Speculative Thinking Yang et al. (2025d)
+SPIRIT Cui et al. (2025)
+ITC Analysis Wang et al. (2025a)
+Think when needed Yang et al. (2025c)
+THINKPRUNE Hou et al. (2025)
+TALE Han et al. (2024)
+TokenSkip Xia et al. (2025)
+TOPS Yang et al. (2025e)
+efficient reasoning Arora and Zanette (2025)
+TWT Xu et al. (2025a)
+Z1 Yu et al. (2025b)
+Count 28 24 20 11 8 6 6 5 5 5
+Table 7: A total of 40 studies on LRM efficiency before May 2025 were included. Benchmarks that appeared more than four times are listed.
 cate. subcatrgories Explanation and cases
@@
 Table 9: The subcategory descriptions and cases of instruction following questions and analysis questions.
+Prompt for construction workflow for S1-Bench
+Data Generation Prompt
+Generate 50 pairs of questions and answers in both Chinese and English based on the category’s name, definition, and specific simplicity criteria. The following conditions must be satisfied:
+1. Questions must be naturally and clearly expressed, unambiguous, and free of intentional traps.
+2. Answers must be unique or easily falsifiable, with no possibility of multiple correct answers.
+3. Make the questions as diverse as possible.
+# Category Name and Definition:
+{name_and_definition}
+# Specific Simplicity Criteria:
+{criteria}
+# Cases:
+## English question:
+{question_en}
+## English Answer:
+{answer_en}
+## Chinese question:
+{question_zh}
+## Chinese Answer:
+{answer_zh}
+Please generate 50 pairs of Chinese and English questions and answers in the following format: [question]English-question[answer]English-answer[question]Chinese-question[answer]Chinese-answer…
+Start generating:
+Quality Discrimination Prompt
+Given a question, its answer, and its category, please analyze from the following perspectives as comprehensively as possible:
+1. Whether the question belongs to the specified category and meet the Specific Simplicity Criteria.
+2. Whether the question is easy, clear, unambiguous, and has an absolutely unique answer.
+3. Whether the answer is absolutely correct; if not, what the correct answer should be.
+4. Whether the question is similar to other given questions, and if similar, whether more diverse questions can be generated.
+# Category Name and Definition:
+{name_and_definition}
+# Specific Simplicity Criteria:
+{criteria}
+# Question and Answer:
+{question_with_answer}
+# Other Questions:
+{questions_list}
+Begin your analysis, aiming to be as detailed and comprehensive as possible:
+Difficulty Reduction Prompt
+Given a question and answer that are too complex for the model to answer correctly, you need to further reduce their difficulty while trying to:
+- Ensure the question aligns with the Category Name and Definition.
+- Ensure the question meets the Specific Simplicity Criteria.
+# Category Name and Definition:
+{name_and_definition}
+# Specific Simplicity Criteria:
+{criteria}
+# Question and Answer:
+{question_with_answer}
+The new question and answer:
+Table 10: “Category Name and Definition” refers to the subcategory name and its definition, while Specific Simplicity Criteria refers to the simplicity requirements specific to the main category.
+Model Model ID URL
+Qwen2.5-7B Qwen2.5-7B-Instruct https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
+Llama3.1-8B Llama-3.1-8B-Instruct https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
+Mistral-8B Ministral-8B-Instruct-2410 https://huggingface.co/mistralai/Ministral-8B-Instruct-2410
+Gemma2-9B gemma-2-9b-it https://huggingface.co/google/gemma-2-9b-it
+Qwen2.5-14B Qwen2.5-14B-Instruct https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
+Qwen2.5-32B Qwen2.5-32B-Instruct https://huggingface.co/Qwen/Qwen2.5-32B-Instruct
+Qwen2.5-72B Qwen2.5-72B-Instruct https://huggingface.co/Qwen/Qwen2.5-72B-Instruct
+Llama3.3-70B Llama-3.3-70B-Instruct https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct
+DeepSeek-v3 DeepSeek-V3-0324 https://huggingface.co/deepseek-ai/DeepSeek-V3-0324
+Table 11: Mapping of LLM abbreviations and IDs used in this paper, with their open-source URLs.
+Model ID Abbreviation Base Model Alg. Size
+DeepSeek
+DeepSeek-R1-Distill-Qwen-1.5B DS-R1-1.5B Qwen2.5-Math-1.5B SFT 800K
+DeepSeek-R1-Distill-Qwen-7B DS-R1-7B Qwen2.5-Math-7B SFT 800K
+DeepSeek-R1-Distill-Llama-8B DS-R1-8B Llama-3.1-8B SFT 800K
+DeepSeek-R1-Distill-Qwen-14B DS-R1-14B Qwen2.5-14B SFT 800K
+DeepSeek-R1-Distill-Qwen-32B DS-R1-32B Qwen2.5-32B SFT 800K
+DeepSeek-R1-Distill-Llama-70B DS-R1-70B Llama-3.3-70B-Instruct SFT 800K
+DeepSeek-R1 DS-R1 DeepSeek-V3-0324 SFT&RL 800K&–
+Qwen
+QwQ-32B QwQ-32B Qwen2.5-32B
+Qwen3-235B-A22B Qwen3-A22B Qwen3-235B-A22B-Base SFT&RL –&–
+Qwen3-30B-A3B Qwen3-A3B Qwen3-30B-A3B-Base SFT&RL –&–
+Qwen3-32B Qwen3-32B Qwen3-32B-Base SFT&RL –&–
+Qwen3-14B Qwen3-14B Qwen3-14B-Base SFT&RL –&–
+Qwen3-8B Qwen3-8B Qwen3-8B-Base SFT&RL –&–
+Qwen3-1.7B Qwen3-1.7B Qwen3-1.7B-Base SFT&RL –&–
+qihoo360
+Light-R1-7B-DS L-R1-7B-DS DeepSeek-R1-Distill-Qwen-7B SFT 3K
+Light-R1-14B-DS L-R1-14B-DS DeepSeek-R1-Distill-Qwen-14B SFT&RL 3K&–
+Light-R1-32B-DS L-R1-32B-DS DeepSeek-R1-Distill-Qwen-32B SFT 3K
+Light-R1-32B L-R1-32B Qwen2.5-32B-Instruct SFT&DPO 73K&–
+simplescaling
+s1.1-7B s1.1-7B Qwen2.5-7B-Instruct SFT 1K
+s1.1-14B s1.1-14B Qwen2.5-14B-Instruct SFT 1K
+s1.1-32B s1.1-32B Qwen2.5-32B-Instruct SFT 1K
+LG AI Research
+EXAONE-Deep-2.4B EXAONE-2.4B EXAONE-3.5-2.4B-Instruct SFT&DPO&RL 1.6M&20K&10K
+EXAONE-Deep-7.8B EXAONE-7.8B EXAONE-3.5-7.8B-Instruct SFT&DPO&RL 1.6M&20K&10K
+EXAONE-Deep-32B EXAONE-32B EXAONE-3.5-32B-Instruct SFT&DPO&RL 1.6M&20K&10K
+NVIDIA
+Llama-3.1-Nemotron-Nano-8B-v1 Nemotron-8B Llama-3.1-8B-Instruct SFT&RL –&–
+Llama-3.3-Nemotron-Super-49B-v1 Nemotron-49B Llama-3.3-70B-Instruct SFT&RL –&–
+NovaSky
+Sky-T1-32B-Flash Sky-T1-32B Qwen2.5-32B-Instruct SFT&SimPO 17K&10K
+Table 12: The open-source LRMs details evaluated for S1-Bench.
+Prompt for Correctness Evaluation
+Evaluation on S1-Bench
+**Question:**
+{question}
+**Model Answer:**
+{model_answer}
+**Ground Truth:**
+{ground_truth}
+Your task is to evaluate whether the model’s answer is correct.
+An answer is considered correct as long as it contains the ground truth (regardless of how complex or detailed the description is).
+If there are parenthetical notes after the ground truth, then there may be multiple correct answers.
+In this case, the given answer is just one example, and any answer that meets the requirements specified in the notes can be considered correct.
+Additionally, some reasonably uncertain supplementary information is also considered appropriate, including more details, possibilities, and expanded discussion.
+You should focus more on whether the reply contains the correct answer.
+You need to output a standard JSON, providing your explanation of the evaluation in the "explain" field, and giving the evaluation result in the "result" field, where 1 means the answer is correct and 0 means it is incorrect.
+Your action should follow the given format: "explain": "", "result": 0/1
+Table 13: Prompt for Correctness Evaluation.
+Model Standard Readable but Malformed Unreadable
+100 101 200 201 202 203 204 205 206 207 300 301
+Qwen3-A22B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Qwen3-A3B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+QwQ-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Qwen3-32B 99.91 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.09
+Qwen3-14B 99.95 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.05
+Qwen3-8B 99.95 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.05
+Qwen3-1.7B 99.81 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.19
+Hunyuan-T1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+DS-R1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+DS-R1-70B 99.91 0.00 0.09 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+DS-R1-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+DS-R1-14B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+DS-R1-8B 99.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24 0.00 0.24
+DS-R1-7B 99.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.76
+DS-R1-1.5B 97.58 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.42
+Sky-T1-32B 95.26 0.00 0.62 0.09 0.19 0.00 0.28 0.00 0.00 3.03 0.00 0.52
+Nemotron-49B 66.07 33.93 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Nemotron-8B 58.06 26.26 0.00 0.00 0.00 0.09 0.00 0.00 0.00 15.02 0.00 0.57
+L-R1-32B 95.07 0.00 0.00 0.00 0.00 0.00 0.81 0.00 0.00 3.03 0.00 1.09
+L-R1-32B-DS 99.81 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.19
+L-R1-14B-DS 99.19 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.81
+L-R1-7B-DS 99.67 0.00 0.05 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
+s1.1-32B 99.53 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.43
+s1.1-14B 97.39 0.00 0.00 0.14 0.00 0.00 0.24 0.00 0.00 0.00 0.00 2.23
+s1.1-7B 88.96 0.00 0.00 7.96 0.09 0.00 0.00 0.00 0.00 0.09 0.00 2.89
+EXAONE-32B 67.39 32.42 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.19
+EXAONE-7.8B 65.83 32.23 0.00 0.00 0.05 0.47 0.00 0.00 0.00 0.14 0.00 1.28
+EXAONE-2.4B 81.42 15.83 0.00 0.09 0.00 0.05 0.00 0.00 0.00 0.05 0.00 2.56
+Table 14: Format type rates under top-p sampling.
+Model Standard Readable but Malformed Unreadable
+100 101 200 201 202 203 204 205 206 207 300 301
+Qwen3-A22B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Qwen3-A3B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+QwQ-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Qwen3-32B 99.76 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
+Qwen3-14B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Qwen3-8B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Qwen3-1.7B 99.76 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
+Hunyuan-T1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+DS-R1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+DS-R1-70B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+DS-R1-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+DS-R1-14B 99.76 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
+DS-R1-8B 99.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24 0.00 0.24
+DS-R1-7B 97.87 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.13
+DS-R1-1.5B 91.94 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 8.06
+Sky-T1-32B 99.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.47 0.00 0.24
+Nemotron-49B 60.90 39.10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+Nemotron-8B 55.21 26.78 0.00 0.00 0.00 0.00 0.00 0.00 0.00 16.35 0.00 1.66
+L-R1-32B 85.55 0.24 0.00 0.24 0.71 0.24 0.95 0.00 0.00 6.64 2.61 2.84
+L-R1-32B-DS 99.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.71
+L-R1-14B-DS 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
+L-R1-7B-DS 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
+s1.1-32B 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
+s1.1-14B 95.97 0.00 0.00 0.24 0.00 0.00 0.24 0.00 0.00 0.00 0.00 3.55
+s1.1-7B 87.91 0.00 0.00 6.64 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.45
+EXAONE-32B 65.88 33.89 0.00 0.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
+EXAONE-7.8B 63.51 33.65 0.00 0.00 0.00 0.24 0.00 0.00 0.00 0.24 0.00 2.37
+EXAONE-2.4B 78.91 15.88 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.21
+Table 15: Format type rates under greedy decoding setting.
+Model Size acc (Loose) acc (Strict) L-Corr $\uparrow$ S-Corr $\uparrow$ Tokens $\downarrow$
+Qwen3-235B-A22B 235B 100.00 100.00 100.00 100.00 702.70
+Qwen3-30B-A3B 30B 100.00 100.00 100.00 100.00 636.35
+QwQ-32B 32B 100.00 100.00 100.00 100.00 750.41
+Qwen3-32B 32B 99.76 99.76 99.76 99.76 673.62
+Qwen3-14B 14B 99.76 99.76 100.00 100.00 597.06
+Qwen3-8B 8B 99.76 99.76 100.00 100.00 649.45
+Qwen3-1.7B 1.7B 99.53 99.53 99.76 99.76 579.01
+Hunyuan-T1 100.00 100.00 100.00 100.00 541.09
+DS-R1 671B 100.00 100.00 100.00 100.00 621.89
+DS-R1-70B 70B 99.76 99.76 100.00 100.00 469.78
+DS-R1-32B 32B 100.00 100.00 100.00 100.00 428.46
+DS-R1-14B 14B 99.29 99.29 99.76 99.76 463.52
+DS-R1-8B 8B 97.63 97.39 99.76 99.53 452.11
+DS-R1-7B 7B 94.31 94.31 97.87 97.87 436.87
+DS-R1-1.5B 1.5B 76.54 76.54 91.94 91.94 473.67
+Sky-T1-32B 32B 99.53 99.05 99.76 99.29 157.12
+Nemotron-49B 49B 99.53 99.53 100.00 100.00 337.94
+Nemotron-8B 8B 84.60 77.73 98.34 81.99 446.62
+L-R1-32B 32B 92.18 85.78 94.55 85.78 996.36
+L-R1-32B-DS 32B 99.29 99.29 99.29 99.29 528.45
+L-R1-14B-DS 14B 98.82 98.82 98.82 98.82 664.28
+L-R1-7B-DS 7B 92.65 92.65 98.82 98.82 514.60
+s1.1-32B 32B 98.82 98.82 98.82 98.82 983.38
+s1.1-14B 14B 95.97 95.50 96.45 95.97 786.30
+s1.1-7B 7B 94.31 87.68 94.55 87.91 630.52
+EXAONE-32B 32B 97.63 97.39 100.00 99.76 746.89
+EXAONE-7.8B 7.8B 86.73 86.49 97.63 97.16 947.92
+EXAONE-2.4B 2.4B 72.99 72.99 94.79 94.79 1394.72
+Table 16: Main results in the greedy decoding setting on the S1-Bench, sorted by model family. Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst.
+Figure 9: Average response token counts on the 28 subcategories, which is the average result of five generations under top-p sampling.
+Prompts for Solution Segmentation
+# Task Description:
+Your task is to segment the given Chain of Thought according to the following rules:
+1. Segmentation positions:
+(1) Please identify and extract all sub-reasoning processes from the Chain of Thought that meet the following condition: They explicitly arrive at a conclusion (including cases phrased as questions, e.g., "right?") that is directly consistent with the Ground Truth.
+Reasoning processes that only indirectly support the Ground Truth or result in partially aligned conclusions should be excluded.
+(2) After clearly reaching the conclusion consistent with the Ground Truth, insert the segmentation marker <split> to distinctly separate each qualifying sub-reasoning process.
+2. Output Restriction:
+(1) You should only directly output the segmentation result without adding any additional supplements.
+(2)Except for inserting the <split> separator, you must not make any other modifications to the original Chain of Thought, not even minor character-level changes such as punctuation, spacing, or capitalization.
+In other words, after removing all <split> separators, the output text must match the original Chain of Thought exactly.
+# Some cases:
+## Case 1:
+### Question:
+What is the value of x if x plus 3 equals 8?
+### Ground Truth:
+5
+### Chain of Thought:
+Okay, so I need to figure out what x is if x plus 3 equals 8. The problem says "x plus 3 equals 8," so in mathematical terms, that would be: x + 3 = 8
+Alright, now I need to solve for x.
+From what I remember in algebra, the goal is to isolate x on one side of the equation.
+That means I need to get rid of that "+ 3" that’s being added to x.
+To do that, I should perform the opposite operation.
+Since it’s adding 3, the opposite would be subtracting 3.
+So, if I subtract 3 from both sides, that should keep things equal. Let me try that: x + 3 - 3 = 8 - 3
+Simplifying both sides, on the left side, the +3 and -3 cancel each other out, leaving just x. And on the right side, 8 minus 3 is 5. So that simplifies to: x = 5
+Wait, let me check if that makes sense. If I substitute x back into the original equation, does it hold true? Let’s see:
+Original equation: x + 3 = 8
+Substitute x = 5: 5 + 3 = 8
+5 plus 3 is indeed 8, so that works out.
+Another way might be to consider what number, when you add 3 to it, gives 8.
+So starting from 8, if I take away 3, I should get back to the original number.
+That’s the same as subtracting 3 from 8, which again gives 5.
+So that’s consistent with the first method.
+Wait a second, what if the original problem was written differently? The problem says "x plus 3 equals 8." In symbols, that’s unambiguous: x + 3 = 8. So no, I think that’s right.
+*Final Answer**
+The value of x is 555.
+### Output:
+Okay, so I need to figure out what x is if x plus 3 equals 8. The problem says "x plus 3 equals 8," so in mathematical terms, that would be:x + 3 = 8
+Alright, now I need to solve for x.
+From what I remember in algebra, the goal is to isolate x on one side of the equation.
+That means I need to get rid of that "+ 3" that’s being added to x.
+To do that, I should perform the opposite operation.
+Since it’s adding 3, the opposite would be subtracting 3.
+So, if I subtract 3 from both sides, that should keep things equal. Let me try that: x + 3 - 3 = 8 - 3
+Simplifying both sides, on the left side, the +3 and -3 cancel each other out, leaving just x. And on the right side, 8 minus 3 is 5. So that simplifies to: x = 5<split>
+Wait, let me check if that makes sense. If I substitute x back into the original equation, does it hold true? Let’s see:
+Original equation: x + 3 = 8
+Substitute x = 5: 5 + 3 = 8
+5 plus 3 is indeed 8, so that works out.<split>
+Another way might be to consider what number, when you add 3 to it, gives 8.
+So starting from 8, if I take away 3, I should get back to the original number.
+That’s the same as subtracting 3 from 8, which again gives 5.<split> So that’s consistent with the first method.
+Wait a second, what if the original problem was written differently? The problem says "x plus 3 equals 8." In symbols, that’s unambiguous: x + 3 = 8. So no, I think that’s right.
+*Final Answer**
+The value of x is 555.<split>
+…(Other examples are omitted.)
+# Eval Target:
+## Question:
+{question}
+## Ground Truth:
+{ground_truth}
+## Chain of Thought:
+{thinking_process}
+## Output:
+Table 17: Prompts for Solution Segmentation.
+Figure 10: Maximum similarity between each segment and all preceding segments for LRMs across four categories.
+Prompts for Error Analysis
+Prompts for samples whose final answer is correct
+# Task Description:
+You will receive a Question, its corresponding Ground Truth, and a Chain of Thought(COT) generated by a LLM for that Question.
+Your task is to carefully analyze the CoT and assign it to one of the two predefined categories listed below.
+# Categories:
+1: The CoT ***includes explicit incorrect conclusions*** in intermediate reasoning steps.
+2: The CoT ***doesn’t include any explicit incorrect conclusion*** in intermediate reasoning steps.
+Output your evaluation in the following format:
+# TheReason: [note: Conduct a step-by-step analysis, stating if and where explicit incorrect conclusions occur in the COT.]
+# ErrorType: [note: Summarize each incorrect conclusion into a specific error type using a phrase of less than 5 words, such as factual inaccuracies, logical fallacies, comprehension mistakes, calculation errors, formatting issues, and so forth, to better conduct further evaluation and analysis.
+Directly output a Python list, where each element represents the error type of a specific incorrect conclusion in the CoT.
+If there are no incorrect conclusions, return an empty list.]
+# TheCategory: [note: Provide your classification based on your analysis using only the number "1" or "2". Do not add any additional text.]
+# Question:
+{question}
+# Ground Truth:
+{gound_truth}
+# COT:
+{thinking_process}
+# TheReason:
+Prompts for samples whose final answer is incorrect
+# Task Description:
+You will receive a Question, its corresponding Ground Truth, and a Chain of Thought(COT) generated by a LLM for that Question.
+Your task is to carefully analyze the CoT and assign it to one of the two predefined categories listed below.
+# Categories:
+1: Regardless of whether the CoT ultimately arrives at the correct final answer or not, ***the correct answer is explicitly mentioned at least once*** within the reasoning steps (even if it is not ultimately adopted).
+2: ***The correct answer is never explicitly mentioned or referenced*** at any point within the reasoning steps.
+Output your evaluation in the following format:
+# TheReason: [note: Conduct a step-by-step analysis, explicitly stating whether and where a correct answer is mentioned within the reasoning steps.]
+# TheCategory: [note: Provide your classification based on your analysis using only the number "1" or "2". Do not add any additional text.]
+# Question:
+{question}
+# Ground Truth:
+{answer}
+# COT:
+{thinking_part}
+# TheReason:
+Table 18: Prompts for Error Analysis.
+Prompts for classify the “gut moment” for English questions
+# Overall Task Description
+You will be given the beginning portion of a response written by a large language model when answering a question.
+Your task is to classify the response into one of the following four categories based on the initial comment about **the difficulty of the question**.
+Important: Only consider the **initial comment** on difficulty made in the response.
+If the model later changes its assessment, please ignore those later revisions—focus solely on the first difficulty-related comment.
+You must assign the response to exactly one of the four categories below:
+Category 1: Initial comment indicates the question is easy. Initial comment includes phrases like “simple,” “basic,” “straightforward,” “common,” etc., clearly stating the question is easy.
+Category 2: Initial comment indicates the question is difficult: Initial comment includes phrases like “a bit complex,” “somewhat difficult,” “challenging,” “tricky,” etc., clearly stating the question is hard.
+Category 3: Initial comment is neutral or suggests moderate difficulty: Initial comment includes phrases like “interesting,” “not hard,” “not complex,” “fun,” etc., without a clear indication of difficulty level.
+Category 4: No comment on difficulty: The response contains no mention of difficulty at all.
+Please output your result in the following format: [Explanation of classification]<sep>[Category (just the number)]<sep>[Short phrase containing the comment]
+If it falls into Category 4 (no comment), write "None" for the [Short phrase containing the comment] part.
+# Examples for each category will be shown below.
+## Category 1 Examples:
+Question:
+What is two plus three?
+Response:
+Hmm, the user is asking what two plus three equals.
+This question seems very simple, though it might be a bit confusing for someone just starting to learn math.
+First, I need to figure out what the user’s intent is.
+They might be testing my basic calculation ability or genuinely need help solving the problem.
+Output:
+The initial comment describes the question as "very simple"<sep>1<sep>This question seems very simple
+…(Other examples are omitted.)
+## Category 2 Examples:
+Question:
+What’s the next floor after basement level two?
+Response:
+Hmm, this question seems a bit complex, but I’ll try to understand it first.
+The question is “What’s the next floor after basement level two?” First, I need to clarify what “basement level two” means.
+Usually, the first basement is the deepest, and then going up you have ground floor, first floor, and so on.
+Output:
+The initial comment says it’s “a bit complex”<sep>2<sep>This question seems a bit complex
+…(Other examples are omitted.)
+## Category 3 Examples: Question:
+“Doctors recommend drinking at least eight glasses of water a day.” Is this sentence about health or education?
+Response:
+Hmm, this question seems not complicated, though if you think carefully it might be a bit tricky.
+First, I need to understand the meaning of the sentence “Doctors recommend drinking at least eight glasses of water a day.”
+Output:
+The initial comment says it’s “not complicated,” even though it later mentions it might be tricky. Classification should follow the initial comment<sep>3<sep>This question doesn’t seem complicated
+…(Other examples are omitted.)
+## Category 4 Examples:
+Question:
+All birds have feathers. A sparrow is a bird. Does a sparrow have feathers?
+Response:
+Hmm, the user is asking a question about birds and feathers—specifically, that all birds have feathers, a sparrow is a bird, so does the sparrow have feathers?
+This looks like a logic reasoning question.
+The user might be learning the basics of logic or biology.
+Output:
+Although the response mentions it’s a “logic reasoning question,” it contains no evaluation of difficulty<sep>4<sep>None
+…(Other examples are omitted.)
+# Here is the one you need to classify.
+Question:
+question
+Response:
+response
+Output:
+Table 19: Prompts for classifying the “gut moment” in English questions.
+Categories: instruction following    Model: DeepSeek-R1-Distill-Qwen-7B
+Question:
+Output the number "7" four times, without using separators. Ground Truth:
+$7$ Thinking process:
+To solve the problem of outputting the number "7" four times without using separators, I need to find a way to repeat the digit 7 multiple times without relying on concatenation or other connecting characters.
+One approach is to use mathematical operations that inherently produce the number 7.
+Addition is a simple and effective operation that can be used to achieve this.
+By adding 7 to itself multiple times, I can create the desired sequence of numbers.
+Specifically, adding 7 four times will give me 7 + 7 + 7 + 7, which equals 28.
+This method ensures that the number 7 is used exactly four times without the need for any separators.
+Therefore, the solution is to use the addition operation to repeat the number 7 four times and output the result.
+Final Answer:
+To output the number "7" four times without using separators, you can utilize mathematical operations that inherently produce the number 7 multiple times.
+One straightforward way is to use addition: [7 + 7 + 7 + 7 = 28] This expression shows that the number "7" is used four times in a row without any separators, resulting in the total of 28.
+[7+7+7+777777 + 7 + 7 + 7]
+Table 20: Error Case for LRM.
+Categories: instruction following    Model: hunyuan-T1
+Question:
+{CJK}UTF8gkai输出一个动物的英文单词,并且只有三个字母 Output an English word for an animal, and it must be only three letters long. Ground Truth:
+{CJK}UTF8gkai任何一个由三个字母组成的动物名称都可以(例如:cat) Any animal name consisting of three letters is acceptable (e.g., cat). Thinking process:
+{CJK}UTF8gkai 嗯,用户让我输出一个动物的英文单词,而且必须是三个字母的。首先,我得确定有哪些常见的动物是三个字母的。可能想到的有猫(Cat)、狗(Dog),不过这两个都是三个字母吗?Cat是C-A-T,确实是三个字母,Dog也是D-O-G,三个字母。那这两个应该都可以。不过用户可能需要更多的例子,或者只需要一个?题目里说的是“输出一个”,所以可能只需要一个。 不过等等,可能还有其他选项,比如Ant(蚂蚁)、Ape(猿)、Bat(蝙蝠)、Bear(熊)是四个字母,所以不行。Fish是四个字母,Fox是三个字母,F-O-X,对,Fox也是三个字母。还有Lion是四个字母,不行。Tiger也是五个字母。Rat是三个字母,R-A-T。还有Sheep是五个字母,Cow是三个字母,C-O-W。所以可能的选择还挺多的。 不过用户可能想要的是最常见的,比如Cat、Dog、Fox、Ant、Rat、Cow这些。需要确认这些单词是否都是正确的动物名称。比如Ant是蚂蚁,没错。Fox是狐狸,Rat是老鼠,Cow是奶牛。这些都是正确的。所以可能需要从这些里面选一个。 不过用户可能希望答案只有一个,所以可能需要随机选一个,或者选最典型的。比如Cat和Dog是最常见的宠物,可能更容易被想到。但Fox也是一个不错的选择,可能更有趣一些。或者像Ape,虽然也是三个字母,但可能不如其他常见。 另外,用户有没有可能有其他需求?比如是否区分大小写?题目里说是英文单词,通常首字母不大写,除非是专有名词。所以应该用小写,比如cat、dog。不过输出的时候可能需要大写首字母,但题目没有特别说明,可能保持小写即可。 再检查一下有没有拼写错误,比如Bat是B-A-T,没错。Rat是R-A-T,正确。Cow是C-O-W,没错。所以这些都可以作为候选答案。现在需要从中选一个作为回答。可能用户希望得到多个例子,但题目要求输出一个,所以可能需要选最典型的。比如Cat或者Dog,因为这两个是最常见的宠物,可能更容易被用户想到。不过也有可能用户希望有变化,比如Fox这样的动物,比较有趣。 或者,用户可能是在测试知识库,看看是否知道除了Cat和Dog之外的其他三个字母的动物。比如Ant,虽然蚂蚁很小,但确实是动物。或者Rat,虽然老鼠可能不太受欢迎,但确实属于动物。Cow是家畜,可能也是合理的选择。 所以,综合考虑,可能需要给出多个例子,但题目要求输出一个。这时候可能需要选择其中一个,比如Cat,或者Dog,或者Fox。例如,回答“Cat”或者“Dog”都是正确的。不过为了多样性,可能选Fox,因为不是所有人都会立刻想到它,而Cat和Dog比较基础。或者用户可能希望得到不同的答案,所以可以选其中一个,比如“Fox”作为回答。 Well, the user asked me to output an English word for an animal, and it must be three letters long.
+First, I need to figure out which common animals have three letters.
+I can think of Cat and Dog, but are both of these three letters?
+Cat is C-A-T, and it is indeed three letters.
+Dog is D-O-G, also three letters.
+So both of these should be valid.
+But the user might need more examples, or just one?
+The task says "output one," so it might only require one.
+But wait, there could be other options, like Ant, Ape, Bat, and Bear, but Bear has four letters, so it doesn’t count.
+Fish has four letters, too.
+Fox is three letters, F-O-X, yes, Fox is three letters.
+Lion has four letters, so that doesn’t work.
+Tiger has five letters.
+Rat is three letters, R-A-T.
+Sheep has five letters.
+Cow is three letters, C-O-W.
+So there are quite a few possibilities.
+However, the user might be looking for the most common ones, like Cat, Dog, Fox, Ant, Rat, and Cow.
+I need to check if these words are all correct animal names.
+For example, Ant is an ant, that’s right.
+Fox is a fox, Rat is a rat, Cow is a cow.
+These are all correct.
+So we might need to choose one from these.
+But the user may want only one answer, so I might need to pick randomly or choose the most typical one.
+For instance, Cat and Dog are the most common pets, so they might be easier to think of.
+But Fox is also a good choice and might be more interesting.
+Or something like Ape, though it’s three letters, might not be as common as the others.
+Also, is there any other requirement from the user?
+For instance, should I differentiate between uppercase and lowercase?
+The task says it’s an English word, and usually, the first letter isn’t capitalized unless it’s a proper noun.
+So, it should be in lowercase, like cat, dog.
+However, when outputting, the first letter might need to be capitalized, but the task doesn’t specify this, so lowercase should be fine.
+Let’s double-check for spelling errors: Bat is B-A-T, correct.
+Rat is R-A-T, correct.
+Cow is C-O-W, correct.
+So these can all be candidate answers.
+Now, I need to choose one as the response.
+The user might want several examples, but the task asks for only one, so I might need to pick the most typical one.
+For example, Cat or Dog, since these are the most common pets and might be easier for the user to think of.
+But Fox is also a good choice, as it’s more interesting.
+Or the user might be testing the knowledge base to see if it knows other three-letter animals besides Cat and Dog.
+For example, Ant, although ants are small, they are indeed animals.
+Or Rat, although rats might not be as popular, they are still animals.
+Cow is livestock, so that could also be a reasonable choice.
+So, taking everything into account, I might need to give multiple examples, but the task requires one.
+In that case, I could choose one, like Cat, or Dog, or Fox.
+For instance, answering “Cat” or “Dog” would both be correct.
+However, for diversity, I might pick Fox, because not everyone will immediately think of it, whereas Cat and Dog are more basic.
+Or the user might want a different answer, so I could choose one of them, like “Fox.” Final Answer:
+**Wolf**
+Table 21: Error Case for LRM.
+Categories: knowledge question    Model: DeepSeek-R1-Distill-Qwen-7B
+Question:
+In which city is the Sydney Opera House located? Ground Truth:
+Sydney. Thinking process:
+Okay, so I need to figure out which city the Sydney Opera House is in.
+I know that Australia has a lot of cities, but I’m pretty sure the Opera House isn’t in Sydney.
+Wait, isn’t it in a different place?
+I think I’ve heard of a place called Petersham before, maybe that’s where it’s located.
+I remember seeing pictures of it in the news once, and they mentioned Petersham as the area where it’s situated.
+But I’m not entirely sure if it’s just a suburb or a major city.
+Maybe I should think about other major cities in Australia.
+There’s Melbourne, which is known as the "Melbourne" city, and then there’s Brisbane, which is more on the northern side.
+Then there’s Sydney, which is the capital, but I don’t think the Opera House is there.
+I think it’s in a regional area instead of the capital city.
+So, putting it all together, I’m pretty confident that the Sydney Opera House is located in Petersham, which is a suburb in New South Wales, Australia.
+But I should double-check to make sure I’m not mixing up the names of places or cities.
+Maybe I can recall any famous landmarks in Petersham.
+I think there’s a big shopping center there called the Petersham Shopping Centre, and I’ve heard that the Opera House is nearby.
+Yeah, that makes sense.
+So, I’m going to go with Petersham as the location.
+Final Answer:
+The Sydney Opera House is located in Petersham, a suburb in New South Wales, Australia.
+Table 22: Error Case for LRM.
+Categories: knowledge question    Model: DeepSeek-R1-Distill-Qwen-7B
+Question:
+Who composed *Symphony No. 9 in D minor, Op. 125*? Ground Truth:
+Ludwig van Beethoven Thinking process:
+Alright, so I need to figure out who composed Symphony No.
+9 in D minor, Op.
+125.
+Let me start by recalling what I know about classical music composers.
+I know that there are several famous composers, but I’m trying to remember if it’s someone like Beethoven, Mozart, or perhaps someone from a later era.
+Wait, Beethoven wrote many symphonies, but I think he didn’t write Symphony No.
+9 in D minor.
+Let me think.
+Beethoven’s 9th symphony is actually in E flat minor, Op.
+67, right?
+So that’s not it.
+Then there’s Beethoven’s Moonlight Sonata, which is different.
+Mozart composed a lot of piano concertos, but I don’t think he wrote a symphony in D minor.
+His compositions are usually in various key signatures, but D minor isn’t one I associate with him.
+Then there’s Schubert, who wrote the famous symphonies, but I don’t recall him composing a symphony in D minor either.
+Wait, maybe it’s someone else.
+I’ve heard of Igor Stravinsky before, but he was more known for his operas like The Rite of Spring.
+Then there’s Paul Simon, but he’s a contemporary composer, so probably not.
+Oh, maybe it’s someone from the classical era.
+Let me think about other composers.
+There’s Shostakovich, but I don’t think he wrote that.
+What about someone like Tchaikovsky?
+No, he wrote symphonies in C minor and F major, but not D minor.
+Hmm, maybe it’s someone more obscure.
+I think I’ve heard of a composer named Igor Kovalsky.
+Wait, was that the composer of this symphony?
+I’m not entirely sure, but I think he’s associated with writing complex symphonies.
+Alternatively, could it be someone like Shoenfield?
+I’m not certain.
+Wait, I think I remember that Igor Kovalsky wrote this symphony.
+It’s sometimes referred to as "The New World" because of its emotional depth and complexity.
+The symphony is in D minor, Op.
+125, and it’s known for its intense, dramatic sound.
+I think that’s correct, but I should double-check to make sure it’s not another composer.
+Another thought: maybe it’s someone from the 19th century.
+I’m trying to recall if any composers from that era wrote such a symphony.
+I’m not sure, but I think Igor Kovalsky is the one.
+He was an American composer, so that makes sense as a modern touch in classical music.
+So, putting it all together, I believe the composer is Igor Kovalsky, and the symphony is known for its dramatic and intense sound.
+I think that’s the correct answer, but I should confirm if there’s any other possibility I might be forgetting.
+Final Answer:
+The composer of Symphony No. 9 in D minor, Op. 125, is Igor Kovalsky. This symphony is often recognized for its dramatic and intense sound, sometimes referred to as "The New World."
+Table 23: Error Case for LRM.
+Categories: reasoning question    Model: s1.1-7B
+Question:
+{CJK}UTF8gkai1.5加3.5等于几? What is 1.5 plus 3.5? Ground Truth:
+5 Thinking process:
+{CJK}UTF8gkai好的,我需要计算1.5加3.5等于多少。首先,这两个数都是小数,处理起来相对简单。我可以在心里快速算一下。 首先,1.5加上3.5。可以将它们看成是1 + 0.5 + 3 + 0.5。这样就可以先将整数部分相加,再将小数部分相加。1+3=4,0.5+0.5=1。然后把两部分加起来是4+1=5。所以答案是5。 另一种方法是将它们视为货币金额。比如,1.5美元加3.5美元。1.5美元加上3美元是4.5美元,再加上0.5美元就是5美元。这也得到同样的结果。 当然,也可以用电子计算器来验证,但我认为手动计算更快。不过,为了确保准确,我再检查一遍。1.5加3.5。10加35等于45,然后小数点一位,所以是4.5。对的,1.5+3.5=4.5。或者用竖式计算也可以: 1.5 +3.5 - - - - - 4.5 对,这样一看也确实是4.5。所以答案应该是4.5。我之前得到的结果是正确的。 Alright, I need to calculate what 1.5 plus 3.5 equals.
+First, both of these numbers are decimals, so it’s relatively simple to handle.
+I can quickly calculate it in my head.
+First, 1.5 plus 3.5.
+You can think of them as 1 + 0.5 + 3 + 0.5.
+This way, I can add the integer parts first, then the decimal parts.
+1 + 3 = 4, and 0.5 + 0.5 = 1.
+Then, adding the two parts together gives 4 + 1 = 5.
+So, the answer is 5.
+Another way is to treat them as currency amounts.
+For example, 1.5 dollars plus 3.5 dollars.
+1.5 dollars plus 3 dollars equals 4.5 dollars, then adding 0.5 dollars gives 5 dollars.
+This also leads to the same result.
+Of course, you could use a calculator to verify, but I think doing it manually is quicker.
+However, to ensure accuracy, I will check again.
+1.5 plus 3.5.
+10 plus 35 equals 45, then move the decimal point one place, so it’s 4.5.
+Yes, 1.5 + 3.5 = 4.5.
+Or you could use column addition: 1.5 +3.5 - - - - - 4.5 Yes, looking at it this way, it indeed adds up to 4.5.
+So the answer should be 4.5.
+The result I got earlier is correct.
+Final Answer:
+{CJK}UTF8gkai答案是 4.54.54.5 The answer is 4.54.54.5
+Table 24: Error Case for LRM.
+Categories: analysis question    Model: DeepSeek-R1-Distill-Qwen-1.5B
+Question:
+In the sentence "Napoleon died in 1821", is "1821" a time or a place name? Ground Truth:
+Time. Thinking process:
+Okay, so I have this sentence: "Napoleon died in 1821." I need to figure out whether "1821" is a time or a place name.
+Hmm, let me think about this step by step.
+First, I know that "Napoleon" is a person, specifically a historical figure.
+So, if it’s the name of a person, then "1821" must be related to their life.
+I’m pretty sure that "1821" is a year.
+I remember that Napoleon was born in 1769 and died in 1821.
+So, 1821 is definitely a year.
+Wait, but what if I’m not sure?
+Let me consider other possibilities.
+Sometimes names can be place names or time periods.
+For example, if someone’s name were a city or a country, the number could be a place.
+But in this case, "Napoleon" is a person, so it’s more likely a year.
+Also, thinking about the sentence structure, it’s written as "Napoleon died in 1821." The structure "in" suggests that 1821 is the place.
+So, in this context, "in" is an article that indicates the location, which would be a place name.
+That makes me think that "1821" is indeed a place name, the year.
+I don’t think it’s a time because time is usually expressed as something like "in the 19th century" or "2023." So, 1821 is specifically a year, not a time period.
+It’s more precise because it’s a specific year when Napoleon died.
+So, putting it all together, "Napoleon" is a person, and "in 1821" is indicating the year he died.
+Therefore, "1821" is a place name, specifically a year.
+Final Answer:
+In the sentence "Napoleon died in 1821," the number "1821" is a place name, specifically a year.
+It is used to indicate the year Napoleon died, which is 1821.
+The use of "in" confirms that it refers to a place, not a time period.
+Therefore, "1821" is a year.
+Table 25: Error Case for LRM.